### v1

# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

Abstract
Spiking Neural Networks (SNNs), as a biologically plausible alternative to Artificial Neural Networks (ANNs), have demonstrated advantages in terms of energy efficiency, temporal processing, and biological plausibility. However, SNNs are highly sensitive to distribution shifts, which can significantly degrade their performance in real-world scenarios. Traditional test-time adaptation (TTA) methods designed for ANNs often fail to address the unique computational dynamics of SNNs, such as sparsity and temporal spiking behavior. To address these challenges, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method specifically designed for SNNs. SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data. We evaluate SPACE on multiple datasets, including CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C and DVS Gesture-C. Furthermore, SPACE demonstrates strong generalization across different model architectures, achieving consistent performance improvements on both VGG9 and ResNet11. Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
1 Introduction
Recent advancements in neuroscience-inspired computing have placed Spiking Neural Networks (SNNs) at the forefront of research as a biologically plausible alternative to traditional Artificial Neural Networks (ANNs). While ANNs have achieved remarkable success in various domains, their reliance on dense, black-box architectures often limits their interpretability and energy efficiency. In contrast, SNNs emulate the sparse, event-driven dynamics of biological neurons, providing several advantages in terms of computational efficiency, temporal processing, and explainability [4, 12, 35]. However, the unique features that make SNNs advantageous—such as their reliance on temporal information and sparse spiking—also pose significant challenges in dynamic real-world environments. SNNs encode information dynamically across time, making them highly sensitive to changes in input data distribution [2]. This sensitivity can result in drastic changes to spiking behavior across the network, ultimately degrading performance when faced with distribution shifts. Understanding and enhancing the robustness of SNNs under such conditions is therefore a critical research direction.
In real-world applications, it is common that the distribution of the training data (a.k.a. source domain) and the test data (a.k.a. target domain) are different due to variations in lighting, background, sensor noise, or environmental conditions. This phenomenon, known as domain shift, can significantly degrade the performance of machine learning models [15, 23]. SNNs are particularly vulnerable to domain shifts, as even minor changes in input data can disrupt the temporal dynamics and spiking behavior in intermediate layers, leading to poor generalization. Traditional test-time adaptation (TTA) solutions such as retraining or fine-tuning on target domain data with help of source data related proxy task [38], statistics of source data [32], or a batch of test data [39], are often impractical in real-world scenarios due to the unavailability of source data or other test samples, particularly in privacy-sensitive or resource-constrained environments. Source-free and single-instance TTA methods [9, 21, 43] offer a promising alternative by enabling the model to adapt to the target domain during inference without access to source data, which is typically defined (and also in this work) as one iteration of training adaptation with one test sample.
Among existing source-free and single-instance TTA methods, MEMO [43] and SITA [21] are two representative approaches, but both face fundamental limitations when applied to SNNs. MEMO optimizes the entropy of the averaged prediction across multiple augmented views of a test sample, enforcing output probability consistency. However, this strategy may not be optimal for SNNs. In SNNs, output probability distributions reflect only high-level semantic information and fail to capture the fine-grained spiking dynamics that encode essential features which prevents MEMO from effectively mitigating domain shifts within SNNs’ internal representations. On the other hand, SITA adapts batch normalization (BN) statistics by shifting them toward those computed from test-time augmentations, but this approach is also unsuitable for SNNs. In SNNs, many models lack BN layers [26, 42], which limits the applicability of BN-based adaptation methods such as SITA in SNNs. Additionally, due to the sparsity of SNN activations, modifying BN statistics does not significantly alter neuronal spiking patterns, further reducing its ability to facilitate effective adaptation.
In this work, we propose a novel method names SPike-Aware Consistency Enhancement (SPACE), which is the first source-free and single-instance TTA approach specifically designed for SNNs. SPACE performs adaptation using only a single test point without access to source data, making it particularly suitable for real-world scenarios where source data are unavailable or privacy-sensitive. By leveraging the inherent spike dynamics of SNNs, SPACE maximizes the consistency in spike-behavior-based local feature maps across augmented samples of the same input, ensuring robust and efficient adaptation in deep SNNs. Unlike existing TTA methods, which often overlook the unique characteristics of SNNs, SPACE directly exploits spike-based representations for adaptation. Our contributions can be summarized as follows
  1. 1.
    To the best of our knowledge, we are the first TTA method tailored for SNNs, which operates effectively using only a single test sample. This addresses the unique challenges of adapting SNNs in scenarios where source data are unavailable, while maintaining efficiency and scalability.
  2. 2.
    Our method introduces a consistency-driven optimization framework to enhance the similarity of spike-behavior-based local feature maps across augmented samples. By leveraging spike dynamics, this approach ensures robust adaptation in deep SNNs.
  3. 3.
    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [24], Tiny-ImageNet (a modified subset of the ImageNet dataset) [6], and the neuromorphic dataset DVS Gesture [1]. Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on VGG9 [37] and ResNet11 [14]. The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
2 Related Work
2.1 SNN-based Deep Learning
Over the years, significant progress has been made in SNN-based deep learning. Many approaches primarily relied on converting pre-trained ANN models into SNNs, enabling SNNs to inherit the representational power of ANNs while benefiting from spike-based computations [3, 8, 17, 20]. However, such conversion methods often suffer from performance degradation due to differences in activation dynamics. To overcome this, direct training of SNNs using surrogate gradient methods has gained traction, enabling end-to-end optimization of spiking models [7, 11, 29, 31, 42]. Rathi et al. [33] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance. These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [26, 40] and spiking recurrent networks [41, 44], which have demonstrated promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations. Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics. According to our experimental results presented in Sec. 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time. Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
2.2 Test-time Adaptation
Test-time adaptation (TTA) is a rapidly growing area in machine learning that tackles the challenge of adapting pre-trained models to unseen or shifted distributions during inference [19, 30, 38, 39]. Test-time training (TTT) [38] performs online updates to model parameters using supervised proxy task on the test data. However, the dependence on source data makes these methods impractical in scenarios where the source domain is unavailable during deployment due to privacy or storage constraints. To overcome source dependency, source-free TTA methods have been proposed, which perform adaptation using only test-time inputs. TENT [39] utilizes a straightforward yet effective entropy minimization approach to optimize batch normalization parameters during test time, without relying on any proxy task during training. Lee et al. [27] employs pseudo-labeling to adjust model predictions.
Although these approaches eliminate the need for source data, they often assume batch-level adaptation, relying on test-time statistics computed over multiple samples. This assumption limits their applicability in real-world settings where only single test points are available at a time. MEMO [43] addresses this scenario, proposing minimizing the marginal output distribution in the increases of the single test point. The objective in MEMO encourages the model to produce consistent predictions across different augmentations, thus enforcing the invariances encoded in these augmentations, while also maintaining confidence in its predictions. Another TTA method, SITA [21], adapts batch normalization statistics at inference by shifting them toward those computed from a batch of augmented versions of the single test instance.
MEMO has shown that enforcing consistency among augmented samples improves TTA performance. However, in deep SNNs, where information is transmitted through discrete spikes, backpropagation tied to output objectives suffers from temporal information loss, making effective TTA challenging. Moreover, MEMO fails to capture the domain shift-induced variations in the internal behavior of the SNN, such as changes in neuron activations and spike patterns, which are critical for robust adaptation in SNNs. Although SITA avoids backpropagation, it also struggles to accurately model the intrinsic temporal dynamics of SNNs, limiting its ability to align augmented samples effectively. To overcome these issues, we enforce consistency by directly aligning the spike patterns and behaviors of augmented samples within the SNN framework. Unlike BN-dependent methods like SITA, our approach works regardless of the presence of BN layers, ensuring broader compatibility with various SNN architectures.
3 Proposed Method
3.1 Preliminary
SNNs emulate biological neurons by transmitting and encoding information through discrete spikes. Among various neuron models, the Leaky Integrate-and-Fire (LIF) neuron [5] is widely adopted for its balance between biological realism and computational efficiency. Its membrane potential dynamics follow
$\tau_{m}\frac{U(t)}{dt}=-U(t)+RI(t),$ (1)
where $U(t)$ represents the membrane potential of the neuron at time $t$, $\tau_{m}$ is the membrane time constant, $R$ denotes the input resistance and $I(t)$ is the input current received from pre-synaptic neurons or inputs. When $U(t)$ exceeds a predefined threshold $U_{th}$, the neuron emits a spike and its membrane potential is reset to a resting value, typically 0 or $U(t)-U_{th}$. Following [11], Eq. 1 is discretized as
$u^{t}_{i}=(1-\frac{1}{\tau_{m}})u_{i}^{t-1}+\frac{1}{\tau_{m}}\sum_{j}w_{ij}o_{j}^{t}.$ (2)
Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons. Fig. 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
The spiking mechanism in SNNs introduces non-differentiability, making it difficult to apply standard gradient-based optimization methods for training or test-time adaptation. To address this, surrogate gradient methods approximate the non-differentiable spike function with a smooth surrogate during the backward pass. A simple approach is the shifted Heaviside step function, where the gradient of $o_{j}$, the spiking activation function of neuron $j$, with respect to $U$ is defined as
$\frac{\partial o_{j}}{\partial U}\triangleq\begin{cases}1,&\text{if }U\geq U_{th},\\ 0,&\text{if }U<U_{th}.\end{cases}$ (3)
While this is not an exact solution, it is valid because a reset occurs after a spike is generated when $U\geq U_{th}$.
These foundational techniques for SNN training, including LIF dynamics and surrogate gradient approximations, form the basis for our proposed method, which focuses on enhancing TTA by ensuring consistency in spiking behavior across augmented samples.
Figure 1: Illustration of spike and membrane potential dynamics in LIF neurons. When the membrane potential reaches the threshold, it is reset by subtracting the threshold value, triggering the neuron to fire a spike.
3.2 SPike-Aware Consistency Enhancement
Figure 2: Overview of the SPike-Aware Consistency Enhancement (SPACE) framework for TTA in SNNs. The test sample is selected from CIFAR-10-C dataset [15] with Gaussian Noise corruption at level 5. The model here follows the VGG9 [37] architecture. The process involves four main steps: 1) Generate an augmented batch from the single test sample. 2) Pass the augmented batch through the model to obtain local feature maps, represented by the spike counts over the processing time. 3) Adapt the model by maximizing the similarity across the local feature maps of the augmented samples. 4) Use the adapted model to predict the label of the original test sample.
Algorithm Design Overview. The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Fig. 2. We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample. This augmented batch is passed through the model to obtain local feature maps, represented by spike counts over time. Next, the model is adapted by maximizing the similarity across the feature maps of the augmented samples, promoting consistency in feature representations. Finally, the adapted model predicts the label of the original test sample, ensuring robust performance under test-time conditions. This approach enhances the model’s ability to generalize to unseen test samples, particularly in shifted domains. The overall method is presented in Algorithm 1, and the details are introduced below. The code will be made available after acceptance.
Algorithm 1 SPACE for Test-Time Adaptation in SNNs
1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, SNN model with parameters $\theta$, and learning rate $\eta$
2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the adapted model
3:  Step 1: Generate augmented batch
4:  Initialize augmented batch $\mathcal{B}=\emptyset$
5:  for each $a_{k}$ in subset $\{a_{1},\dots,a_{M}\}$ of $\mathcal{A}$ do
6:     Apply augmentation: $\mathbf{x}_{k}=a_{k}(\mathbf{x})$
7:     Add $\mathbf{x}_{k}$ to $\mathcal{B}$
8:  end for
9:  Step 2: Extract Feature Maps
10:  Pass $\mathcal{B}$ through the SNN feature extractor
11:  Obtain feature maps:
12:  =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
13:  Step 3: Compute Similarity
14:  for each pair $(\mathbf{F}(\mathbf{x}_{i}),\mathbf{F}(\mathbf{x}_{j}))$ in $\mathcal{F}$ do
15:     Compute similarity $\bar{\mathcal{S}}(i,j|\mathbf{x})$ using Eq. 4 $\sim$ Eq. 6
16:  end for
17:  Step 4: Compute loss and update model
18:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Eq. 7
19:  Update parameters using Eq. 8
20:  Step 5: Predict via adapted model
21:  Predict via: $y^{*}=argmax_{y}p_{\theta^{*}}(y|\mathbf{x})$
22:  Return: Prediction result $y^{*}$
Spike-Aware Feature Maps. SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data. No specific training process is required, nor do we impose particular constraints on the model. The only assumptions are that $\theta$ can be adjusted and that the model produces intermediate feature maps $\mathbf{F}(\mathbf{x})$, which is differentiable with respect to $\theta$ and can be utilized for further analysis and adaptation. Here, $\mathbf{x}\in\mathcal{X}$ is a single test point that is presented to $M_{\theta}$. It is worth noting that this is a reasonable assumption, supported by a recent advance [34] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware. Aligned with previous work [21, 43], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective. In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples =$\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$. Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts.
Different from MEMO [43], which adapts the model during test time via the conditional output distribution $p_{\theta}(y|\mathbf{x})$ of the model’s last layer, SPACE leverages the dynamic behavior of intermediate layers in the SNN that plays a crucial role in determining the temporal information and semantic representation of the final output. SNNs encode information through temporal dynamics and spike sparsity. These behavioral characteristics are particularly prominent in the intermediate layers and can be represented by feature maps $\mathbf{F}(\mathbf{x})$. Unlike ANNs, these characteristics in SNNs cannot be fully captured by the output probability distribution alone. Another advantage of intermediate feature maps is its highly sensitivity to changes in the input distribution due to the event-driven nature of SNNs. When the input data undergoes a distribution shift, the spiking patterns in the intermediate layers may change significantly, thereby impacting the final output. Optimizing the consistency of spiking behavior in intermediate layers can effectively mitigate feature variations caused by distribution shifts, thereby enhancing the robustness of models.
To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window. As described in Sec. 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$. The temporal spike patterns of the network are captured in the feature map $\mathbf{O(\mathbf{x}_{i})}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps. $\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch. Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network (typically a CNN [25]), depending on whether the final feature map is pooled into a 1$\times$1 matrix. Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
There are several reasons for choosing the total spike counts as the observation metric. First, SNNs typically operate over a time window ranging from 25 to thousands time steps [4, 13, 26, 33, 36], meaning that spike patterns across many time steps may be identical due to the repetitive nature of the spike dynamics. By comparing the cumulative spike counts instead of individual time-step patterns, we can effectively capture the temporal accumulation of spike activity while reducing redundancy. Second, the total spike counts reflects differences between augmented samples based on their overall spike dynamics, which is essential for our goal of aligning feature maps across augmentations. Finally, this representation significantly reduces the computational cost, as it condenses the temporal information into an aggregate value without requiring a step-by-step comparison of spike patterns.
Feature Maps Alignment. The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data. When the feature maps of augmented samples are highly similar, the model becomes less sensitive to variations at test time, thereby improving its ability to generalize to new, unseen, and unlabeled data. To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$. Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution:
$\mathbf{P}_{c}(\mathbf{x}_{i})=\textit{softmax}(\mathbf{F}_{c}(\mathbf{x}_{i})).$ (4)
Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ represent the local feature vectors corresponding to channel $c$ of the augmented sample $\mathbf{x}_{i}$. This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively. To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
$\mathcal{S}_{c}(i,j|\mathbf{x})=\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (5)
$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{S}_{c}(i,j|\mathbf{x}).$ (6)
The similarity measures the degree of overlap between the feature distributions, with higher values indicating greater consistency in spike-based representations across augmentations. Based on this similarity measure, we define the loss function for model adaptation given as
$\mathcal{L}(\theta;\mathbf{x})\triangleq\sum_{1\leq j<i\leq M}\left(1-\bar{\mathcal{S}}(i,j|\mathbf{x})\right).$ (7)
Then we adapt the model via gradient descent for only one iteration given as
$\theta^{*}\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta;\mathbf{x}),$ (8)
where $\eta$ is learning rate here. By minimizing this loss function, we enforce consistency across the feature maps of different augmentations. This reduces the variability induced by augmentations, ensuring that the model learns stable and robust feature representations. Furthermore, by promoting feature consistency, this approach serves as a regularization mechanism, preventing over-fitting to specific augmentations and encouraging the model to focus on more fundamental and generalizable features.
While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure. Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) (see Sec. 6 in Supplementary Material). However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
Finally, we predict the classification via an updated model.
Discussion. In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally. Since channels may have different active neurons, directly comparing the entire feature map fails to account for this variability. Instead, computing the similarity per channel and averaging the results focuses on the stability of feature representations within each channel. This isolates the contributions of active neurons, ensuring the similarity measure reflects meaningful changes rather than noise. It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels. Our experiments in Sec. 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach. Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.
4 Experimental Results
Datasets. The experiments here were conducted on three benchmark datasets: CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C [15], each containing a variety of common image corruptions. CIFAR-10-C includes 18 corruption types, while CIFAR-100-C features 19 types, covering weather, noise, and digital distortions. Both CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes of 32$\times$32 color images, respectively, similar to CIFAR-10 and CIFAR-100 [24]. Tiny-ImageNet-C, derived from the Tiny-ImageNet dataset, contains 15 corruption types and 200 classes of 64$\times$64 color images. In addition, we also considered a neuromorphic dataset, DVS Gesture-C [18], and the related experiments are provided in the Supplementary Material Sec. 7.
Experimental Setup. All experiments were conducted under the highest corruption level (level=5). Following the experimental setup of MEMO [43], we applied AugMix [16] augmentation with a batch size of 32 to each test sample to generate a diverse set of augmented samples. The backbone model for most experiments is an SNN-based VGG architecture with Batch-Normalization Through Time (BNTT) [22], featuring BN layers at each time step, with 25 inference steps. To demonstrate the general applicability of the proposed SPACE method, additional experiments were conducted using an SNN-based ResNet-11 model [26] on the CIFAR-10-C dataset, with inference steps set to 30.
Baseline TTA Methods. We compared our approach with models pre-trained on clean datasets without any test-time adaptation (referred to as w.o. TTA). Moreover, since few TTA methods are specifically designed for SNNs, we selected two most representative traditional TTA methods, MEMO [43] and SITA [21], as our baselines.
4.1 Main Results
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 72.38% 73.06% 77.73% 77.98%
shot 74.70% 74.15% 79.50% 79.34%
speckle 71.75% 71.67% 77.28% 77.38%
impulse 58.57% 58.41% 65.74% 69.41%
blur defocus 63.05% 62.94% 65.61% 71.59%
gaussian 57.03% 57.36% 59.93% 68.31%
motion 64.44% 64.36% 66.24% 72.14%
zoom 71.33% 70.72% 72.61% 74.67%
weather snow 76.32% 76.67% 77.38% 78.43%
fog 43.57% 43.40% 45.47% 52.80%
frost 75.72% 75.24% 78.64% 79.59%
digital brightness 82.44% 82.51% 83.05% 83.22%
contrast 22.54% 22.02% 23.83% 23.85%
elastic_transform 75.01% 74.74% 76.24% 75.49%
pixelate 70.25% 69.62% 73.25% 76.24%
jpeg_compression 84.28% 84.16% 86.05% 82.88%
spatter 78.42% 78.61% 79.75% 77.52%
saturate 80.85% 80.97% 82.03% 82.54%
Average 67.93% 67.81% 70.53% 72.41%
Table 1: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO [43] and SITA [21], which are current state-of-the-art methods for source-free and single-instance TTA. While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation. Similarly, SITA, despite not using backpropagation, fails to capture the intrinsic temporal dynamics of SNNs, rendering it ineffective. To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using the VGG architecture, and on CIFAR-10-C with the ResNet-11 architecture to demonstrate the broad applicability of our proposed method.
4.1.1 Performance Comparison on VGG Model
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 42.51% 43.01% 43.46% 44.71%
shot 45.40% 44.83% 45.41% 46.73%
speckle 43.35% 43.14% 43.53% 44.84%
impulse 25.50% 25.48% 26.07% 27.99%
blur defocus 42.15% 42.23% 42.47% 43.35%
gaussian 37.76% 37.88% 38.29% 39.19%
motion 41.87% 42.11% 42.55% 43.46%
zoom 46.77% 46.93% 47.21% 48.42%
glass 42.84% 43.29% 43.90% 44.47%
weather snow 44.88% 44.49% 44.74% 45.70%
fog 16.51% 16.65% 16.76% 17.62%
frost 44.12% 43.84% 44.34% 44.98%
digital brightness 49.36% 49.53% 49.95% 50.41%
contrast 5.42% 5.56% 5.62% 5.69%
elastic_transform 51.33% 50.99% 51.24% 51.45%
pixelate 52.35% 52.48% 52.82% 53.81%
jpeg_compression 58.47% 58.71% 59.29% 58.96%
spatter 47.76% 47.82% 48.63% 48.95%
saturate 46.91% 47.17% 47.17% 48.22%
Average 41.33% 41.38% 41.76% 42.58%
Table 2: Performance comparison of different methods on CIFAR-100-C dataset at level 5 corruption, using the VGG11 architecture.
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 12.43% 12.71% 13.64% 16.71%
shot 15.20% 15.17% 16.45% 19.44%
impulse 8.74% 8.87% 9.51% 11.72%
defocus 7.58% 7.56% 7.51% 9.62%
blur motion 14.48% 14.57% 14.38% 16.57%
zoom 13.59% 13.94% 13.92% 15.61%
glass 6.50% 6.57% 6.60% 7.31%
weather snow 15.58% 15.72% 15.97% 18.66%
fog 5.27% 5.32% 4.91% 6.01%
frost 18.06% 18.05% 17.83% 21.02%
digital brightness 15.01% 14.82% 14.74% 17.53%
contrast 1.32% 1.38% 1.44% 1.51%
elastic_transform 20.53% 21.45% 21.00% 21.96%
pixelate 31.62% 32.20% 31.39% 31.89%
jpeg_compression 31.15% 31.51% 30.82% 31.63%
Average 14.47% 14.66% 14.67% 16.48%
Table 3: Performance comparison of different methods on Tiny-ImageNet-C dataset at level 5 corruption, using the VGG11 architecture.
Shifted Domain w.o. TTA MEMO SPACE
noise gaussian 72.88% 74.21% 75.90%
shot 74.49% 75.50% 77.81%
speckle 73.12% 74.32% 76.75%
impulse 49.59% 51.35% 56.85%
blur defocus 67.77% 68.11% 68.66%
gaussian 63.57% 64.11% 64.15%
motion 65.14% 65.37% 66.02%
zoom 73.06% 73.96% 73.82%
weather snow 78.11% 78.32% 78.74%
fog 42.84% 43.47% 45.47%
frost 73.12% 73.39% 74.04%
digital brightness 81.71% 81.82% 81.28%
contrast 12.72% 13.63% 14.94%
elastic_transform 74.88% 75.54% 76.79%
pixelate 77.08% 77.54% 80.32%
jpeg_compression 82.47% 83.24% 83.96%
spatter 78.75% 79.46% 80.58%
saturate 79.40% 79.40% 78.70%
Average 67.82% 68.49% 69.71%
Table 4: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the ResNet11 architecture.
The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Tab. 1, Tab. 2, and Tab. 3, demonstrate the effectiveness of SPACE. SPACE outperforms w.o. TTA, MEMO, and SITA in average accuracy across all datasets. On CIFAR-10-C, SPACE achieves 72.41%, improving by 4.48% over w.o. TTA (67.93%), 4.60% over SITA (67.81%), and 1.88% over MEMO (70.53%). On CIFAR-100-C, SPACE reaches 42.58%, surpassing w.o. TTA (41.33%), SITA (41.38%), and MEMO (41.76%). On Tiny-ImageNet-C, SPACE achieves 16.48%, compared to 14.47% for w.o. TTA, 14.66% for SITA, and 14.67% for MEMO. Notably, SPACE consistently improves performance on Tiny-ImageNet-C across all domains, showing its robustness on complex datasets.
Across individual corruption types, SPACE consistently outperforms w.o. TTA, MEMO, and SITA. For instance, on CIFAR-10-C, SPACE shows notable improvements, such as in impulse noise (69.41% for SPACE vs. 58.57% for w.o. TTA, 58.41% for SITA, and 65.74% for MEMO) and fog (52.80% for SPACE vs. 43.57% for w.o. TTA, 43.40% for SITA, and 45.47% for MEMO). On CIFAR-100-C, SPACE similarly excels, including in motion blur (43.46% for SPACE vs. 41.87% for w.o. TTA, 42.11% for SITA, and 42.55% for MEMO).
A particularly striking result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type. For example, in gaussian noise and impulse noise, SPACE leads by a significant margin. Even in simpler corruptions like brightness and pixelate, SPACE outperforms the w.o. TTA, demonstrating its robustness in complex datasets like Tiny-ImageNet-C.
Although SPACE generally outperforms MEMO and SITA, MEMO does slightly better in a few cases, such as jpeg_compression on CIFAR-10-C (86.05% for MEMO vs. 82.88% for SPACE) and on CIFAR-100-C (59.29% for MEMO vs. 58.96% for SPACE). These minor differences suggest MEMO has an edge with specific corruptions, but they do not diminish SPACE’s overall superiority.
4.1.2 Performance Comparison on ResNet Model
We evaluated the SPACE method on the CIFAR-10-C dataset using the ResNet11 architecture, which lacks BN layers, making SITA inapplicable. The results, shown in Tab. 4, demonstrate SPACE’s effectiveness across different model structures. SPACE achieves an average accuracy of 69.71%, improving by 1.22% over MEMO (68.49%) and 1.89% over w.o. TTA (67.82%).
In terms of individual corruption types, SPACE outperforms MEMO and w.o. TTA in most cases. Notable improvements include challenging corruptions such as impulse noise (56.85% vs. 49.59% for w.o. TTA and 51.35% for MEMO) and contrast (14.94% vs. 12.72% for w.o. TTA and 13.63% for MEMO). SPACE also achieves 80.32% on pixelate, outperforming MEMO (77.54%) and w.o. TTA (77.08%). These trends align with those observed in the VGG-based experiments, reaffirming SPACE’s consistency and generalizability.
While MEMO marginally outperforms SPACE on zoom (73.96% vs. 73.82%) and brightness (81.82% vs. 81.28%), these differences are minimal and do not overshadow SPACE’s overall advantages. The results further confirm that SPACE is not dependent on a specific architecture and can adapt effectively to various backbone models.
4.2 Ablation Study
Figure 3: Local Feature Maps Based SPACE Outperforms Global Feature Maps Based SPACE.
Local vs. Global Feature Map Similarity. We compared the performance of local versus global feature map similarity for our proposed method. The results in Fig. 3 clearly illustrate the advantage of using similarity of local feature maps over global feature maps. While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency. As shown in Fig. 3, the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets. This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation. These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
Figure 4: Distribution of similarity values between spike counts based feature maps before and after SPACE adaptation.
Effect of SPACE on Feature Map Similarity. To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Eq. 4 $\sim$ Eq. 6) across the CIFAR-10-C dataset. As shown in Fig. 4, we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Gaussian Blur and Impulse Noise domains. The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process. Specifically, the “After SPACE Adaptation” distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples. This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
5 Conclusion
In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs. SPACE leverages SNNs’ unique spike dynamics to optimize spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods, which often ignore the temporal and sparse nature of SNN computation. Extensive experiments on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C show SPACE achieves consistent performance improvements under severe corruptions, while generalizing effectively to various architectures like VGG9 and ResNet11. These results demonstrate SPACE’s robustness and adaptability to domain shifts, paving the way for its potential use in neuromorphic hardware and energy-efficient AI systems.
Supplementary Material
6 Kernel-Based Consistency Regularization
To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function. The channel-wise feature probability distribution $\mathbf{P}_{c}$ obtained from Eq. 4 can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
$k(\mathbf{z},\mathbf{z}^{\prime})=\text{exp}(-\frac{\|\mathbf{z}-\mathbf{z}^{\prime}\|^{2}}{2\sigma^{2}}),$ (9)
where $\sigma$ is the kernel bandwidth. The kernel function $k(\mathbf{z},\mathbf{z}^{\prime})$ implicitly defines a mapping $\phi$: $\mathcal{Z}\xrightarrow{}\mathcal{H}$, such that the inner product in $\mathcal{H}$ is given by the kernel:
$k(\mathbf{z},\mathbf{z}^{\prime})=\langle\phi(\mathbf{z}),\phi(\mathbf{z}^{\prime})\rangle_{\mathcal{H}}.$ (10)
This allows us to compute relationships between feature vectors in a high-dimensional space without explicitly constructing $\phi(\cdot)$. To align feature distributions from augmented samples, we use the Mean Embedding of Distributions, which maps $\mathbf{P}_{c}(\mathbf{x}_{i})$ into a single point $\mu_{\mathbf{P}_{c}(\mathbf{x}_{i})}$ in RKHS. Specifically, the mean embedding is defined as
$\mu_{\mathbf{P}_{c}(\mathbf{x}_{i})}=\mathbb{E}_{\mathbf{z}_{i}\sim\mathbf{P}_{c}(\mathbf{x}_{i})}[\phi(\mathbf{z}_{i})].$ (11)
Given samples $\mathbf{z}_{i}\sim\mathbf{P}_{c}(\mathbf{x}_{i})$, the embedding can be approximated as
$\mu_{\mathbf{P}_{c}(\mathbf{x}_{i})}\triangleq\frac{1}{D}\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}}).$ (12)
To measure the discrepancy between distributions $\mathbf{P}_{c}(\mathbf{x}_{i})$ and $\mathbf{P}_{c}(\mathbf{x}_{j})$ in RKHS, we employ the Maximum Mean Discrepancy (MMD):
$\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\|\mu_{\mathbf{P}_{c}(\mathbf{x}_{i})}-\mu_{\mathbf{P}_{c}(\mathbf{x}_{j})}\|^{2}_{\mathcal{H}}.$ (13)
Using the kernel trick, this can be computed without explicitly constructing $\phi(\cdot)$:
$\begin{split}\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\frac{1}{D^{2}}\|\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}})-\sum^{D}_{d=1}\phi(\mathbf{z}_{j_{d}})\|^{2}_{\mathcal{H}}.\end{split}$ (14)
Then we average the MMD values across all channels:
$\text{MMD}^{2}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j})).$ (15)
Finally, we integrate the average MMD values into the original loss function $\mathcal{L}$:
$\mathcal{L}^{*}(\theta;\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (16)
Shifted Domain w.o. TTA SPACE SPACE + Kernel Embedding
noise gaussian 72.38% 77.98% 77.85%
shot 74.70% 79.34% 79.34%
speckle 71.75% 77.38% 77.33%
impulse 58.57% 69.41% 69.25%
blur defocus 63.05% 71.59% 71.42%
gaussian 57.03% 68.31% 67.98%
motion 64.44% 72.14% 72.21%
zoom 71.33% 74.67% 74.61%
weather snow 76.32% 78.43% 78.48%
fog 43.57% 52.80% 52.46%
frost 75.72% 79.59% 79.66%
digital brightness 82.44% 83.22% 83.25%
contrast 22.54% 23.85% 23.40%
elastic_transform 75.01% 75.49% 75.42%
pixelate 70.25% 76.24% 76.13%
jpeg_compression 84.28% 82.88% 82.98%
spatter 78.42% 77.52% 77.47%
saturate 80.85% 82.54% 82.64%
Average 67.93% 72.41% 72.41%
Table 5: Performance comparison of SPACE without and with kernel embedding on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
Shifted Domain w.o. TTA MEMO SPACE
DropPixel 44.53% 43.94% 46.97%
DropEvent 50.00% 51.52% 53.41%
RefractoryPeriod 35.16% 34.85% 39.39%
TimeJitter 42.19% 40.91% 42.42%
SpatialJitter 83.98% 82.95% 85.23%
UniformNoise 75.39% 74.24% 75.39%
Average 55.21% 54.74% 57.14%
Table 6: Performance comparison of different methods on DVS Gesture-C dataset at level 5 corruption.
To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using the BNTT model with the VGG9 architecture, as shown in Tab. 5. The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical. This limited effectiveness can be attributed to several factors. First, the distribution differences introduced by augmentations, such as noise, blur, and weather effects, are relatively small, reducing the need for advanced alignment mechanisms like kernel embedding. Additionally, the existing similarity measures in SPACE already effectively capture the relationships between augmented samples, leaving little room for further optimization. Moreover, the added computational complexity of kernel embedding, especially the cost of kernel calculations, may even slightly hinder performance in scenarios where the original approach is already sufficient. These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
7 Performance on DVS Gesture-C Dataset
SNNs are inherently suited for processing event-based neuromorphic data due to their temporal dynamics and energy efficiency. Static image datasets like CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C are commonly used to evaluate robustness to corruption and distribution shifts, but they lack the temporal and asynchronous characteristics of real-world event-based data. DVS-Gesture [1] allows us to evaluate the proposed method’s performance on neuromorphic datasets, emphasizing its compatibility with dynamic vision tasks and data with temporal structure.
Following the state-of-the-art methodology [18], we evaluate the robustness improvements achieved by SPACE using DVS-Gesture-C, a corrupted variant of the standard DVS-Gesture dataset. This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise. These corruptions, implemented via the Tonic API [28], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies. To ensure a comprehensive and stringent assessment, we consistently employ the highest severity level across all experimental evaluations. The pre-trained model used in this study follows a 2CONV-2FC architecture, which was selected for its balance between computational efficiency and representational capacity, ensuring a fair and reliable comparison.
The experimental results demonstrate that the proposed SPACE method achieves superior performance on the DVS-Gesture-C dataset compared to both w.o. TTA and the MEMO method. As shown in Tab. 6, SPACE consistently outperforms w.o. TTA across all six corruption types, with the largest improvements observed for RefractoryPeriod (+4.23%) and DropEvent (+3.41%). Compared to MEMO, SPACE also achieves higher accuracy in most scenarios, with notable gains for RefractoryPeriod (+4.54%) and DropPixel (+3.03%). On average, SPACE achieves an accuracy of 57.14%, significantly surpassing both w.o. TTA (55.21%) and MEMO (54.74%).
While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation. This limitation is evident in the results, where MEMO struggles to handle certain corruption types effectively, such as RefractoryPeriod and TimeJitter. In contrast, SPACE addresses these challenges by fully leveraging the temporal dynamics and event-based nature of SNNs, resulting in more robust and reliable performance across diverse corruption types. These results highlight the robustness and effectiveness of SPACE in improving model performance under challenging conditions, making it a highly suitable method for event-based data with various noise and corruption types.

### v2

@@
 SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data.
-We evaluate SPACE on multiple datasets, including CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C and DVS Gesture-C.
-Furthermore, SPACE demonstrates strong generalization across different model architectures, achieving consistent performance improvements on both VGG9 and ResNet11.
+We evaluate SPACE on multiple datasets.
+Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs (such as VGG and ResNet), Transformer models, and ConvLSTM architectures.
 Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
@@
   3. 3.
-    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [24], Tiny-ImageNet (a modified subset of the ImageNet dataset) [6], and the neuromorphic dataset DVS Gesture [1].
-Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on VGG9 [37] and ResNet11 [14].
+    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [26], Tiny-ImageNet, ImageNet [6], and the neuromorphic dataset DVS Gesture [1].
+Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [24, 40], SNN-ResNet11 [14, 27], Spike-driven Transformer V3 [44], and SNN-ConvLSTM [39].
 The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@
 [34] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance.
-These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [26, 40] and spiking recurrent networks [41, 44], which have demonstrated promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
+These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [27, 43] and spiking recurrent networks [45, 48], which have denoted promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
 A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations.
 Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics.
-According to our experimental results presented in Sec.
-4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time.
+According to our experimental results presented in section 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time.
 Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
@@
 When $U(t)$ exceeds a predefined threshold $U_{th}$, the neuron emits a spike and its membrane potential is reset to a resting value, typically 0 or $U(t)-U_{th}$.
-Following [11], Eq.
-1 is discretized as
+Following [11], Equation 1 is discretized as
 $u^{t}_{i}=(1-\frac{1}{\tau_{m}})u_{i}^{t-1}+\frac{1}{\tau_{m}}\sum_{j}w_{ij}o_{j}^{t}.$ (2)
+Figure 1: Illustration of spike and membrane potential dynamics in LIF neurons.
+When the membrane potential reaches the threshold, it is reset by subtracting the threshold value, triggering the neuron to fire a spike.
 Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons.
-Fig.
-1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
+Figure 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
 The spiking mechanism in SNNs introduces non-differentiability, making it difficult to apply standard gradient-based optimization methods for training or test-time adaptation.
@@
 These foundational techniques for SNN training, including LIF dynamics and surrogate gradient approximations, form the basis for our proposed method, which focuses on enhancing TTA by ensuring consistency in spiking behavior across augmented samples.
-Figure 1: Illustration of spike and membrane potential dynamics in LIF neurons.
-When the membrane potential reaches the threshold, it is reset by subtracting the threshold value, triggering the neuron to fire a spike.
 3.2 SPike-Aware Consistency Enhancement
@@
 The test sample is selected from CIFAR-10-C dataset [15] with Gaussian Noise corruption at level 5.
-The model here follows the VGG9 [37] architecture.
+The model here follows the VGG9 architecture.
 The process involves four main steps: 1) Generate an augmented batch from the single test sample.
@@
 4) Use the adapted model to predict the label of the original test sample.
-Algorithm Design Overview.
-The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Fig.
-2.
+Algorithm Design Overview
+The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Figure 2.
 We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample.
@@
 The code will be made available after acceptance.
-Algorithm 1 SPACE for Test-Time Adaptation in SNNs
-1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, SNN model with parameters $\theta$, and learning rate $\eta$
-2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the adapted model
-3:  Step 1: Generate augmented batch
-4:  Initialize augmented batch $\mathcal{B}=\emptyset$
-5:  for each $a_{k}$ in subset $\{a_{1},\dots,a_{M}\}$ of $\mathcal{A}$ do
-6:     Apply augmentation: $\mathbf{x}_{k}=a_{k}(\mathbf{x})$
-7:     Add $\mathbf{x}_{k}$ to $\mathcal{B}$
-8:  end for
-9:  Step 2: Extract Feature Maps
-10:  Pass $\mathcal{B}$ through the SNN feature extractor
-11:  Obtain feature maps:
-12:  =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
-13:  Step 3: Compute Similarity
-14:  for each pair $(\mathbf{F}(\mathbf{x}_{i}),\mathbf{F}(\mathbf{x}_{j}))$ in $\mathcal{F}$ do
-15:     Compute similarity $\bar{\mathcal{S}}(i,j|\mathbf{x})$ using Eq. 4 $\sim$ Eq. 6
-16:  end for
-17:  Step 4: Compute loss and update model
-18:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Eq. 7
-19:  Update parameters using Eq. 8
-20:  Step 5: Predict via adapted model
-21:  Predict via: $y^{*}=argmax_{y}p_{\theta^{*}}(y|\mathbf{x})$
-22:  Return: Prediction result $y^{*}$
-Spike-Aware Feature Maps.
+Spike-Aware Feature Maps
 SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data.
@@
 To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window.
-As described in Sec.
-3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$.
+As described in subsection 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$.
 The temporal spike patterns of the network are captured in the feature map $\mathbf{O}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps.
 $\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch.
-Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network (typically a CNN [25]), depending on whether the final feature map is pooled into a 1$\times$1 matrix.
-Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
+Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network, depending on whether the final feature map is pooled into a 1$\times$1 matrix.
+Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
 There are several reasons for choosing the total spike counts as the observation metric.
@@
 Finally, this representation significantly reduces the computational cost, as it condenses the temporal information into an aggregate value without requiring a step-by-step comparison of spike patterns.
-Feature Maps Alignment.
+Algorithm 1 SPACE for Test-Time Adaptation in SNNs
+1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, model $\theta$, learning rate $\eta$
+2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the updated model
+3:  Step 1: Generate augmented batch
+4:  Initialize $\mathcal{B}=\emptyset$
+5:  for $a_{k}$ in a subset $\{a_{1},\dots,a_{M}\}$ of $\mathcal{A}$ do
+6:     $\mathbf{x}_{k}=a_{k}(\mathbf{x})$, add $\mathbf{x}_{k}$ to $\mathcal{B}$
+7:  end for
+8:  Step 2: Extract feature maps
+9:  Pass $\mathcal{B}$ through SNN, obtain =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
+10:  Step 3: Compute similarity
+11:  for each pair of ($\mathbf{F}(\mathbf{x}_{i}),\mathbf{F}(\mathbf{x}_{j})$) in $\mathcal{F}$ do
+12:     Compute similarity $\bar{\mathcal{S}}(i,j|\mathbf{x})$ using Equation 4
+13:  end for
+14:  Step 4: Update model parameters
+15:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta$ via SGD with $\eta$
+16:  Step 5: Predict
+17:  $y^{*}=\arg\max_{y}p_{\theta^{*}}(y|\mathbf{x})$
+18:  Return: $y^{*}$
+Feature Maps Alignment
 The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data.
@@
 To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$.
-Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution:
-$\mathbf{P}_{c}(\mathbf{x}_{i})=\textit{softmax}(\mathbf{F}_{c}(\mathbf{x}_{i})).$ (4)
-Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ represent the local feature vectors corresponding to channel $c$ of the augmented sample $\mathbf{x}_{i}$.
+Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ denotes the local feature vectors associated with channel $c$ of the augmented sample $\mathbf{x}_{i}$.
+Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$.
 This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively.
 To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
-$\mathcal{S}_{c}(i,j|\mathbf{x})=\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (5)
-$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{S}_{c}(i,j|\mathbf{x}).$ (6)
+$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
 The similarity measures the degree of overlap between the feature distributions, with higher values indicating greater consistency in spike-based representations across augmentations.
@@
 $\mathcal{L}(\theta;\mathbf{x})\triangleq\sum_{1\leq j<i\leq M}\left(1-\bar{\mathcal{S}}(i,j|\mathbf{x})\right).$ (5)
-Then we adapt the model via gradient descent for only one iteration given as
-$\theta^{*}\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta;\mathbf{x}),$ (8)
-where $\eta$ is learning rate here.
+Then, we adapt the model using stochastic gradient descent (SGD) for a single iteration with the learning rate $\eta$.
 By minimizing this loss function, we enforce consistency across the feature maps of different augmentations.
@@
 While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure.
-Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) (see Sec.
-6 in Supplementary Material).
+Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) in Appendix).
 However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
 Finally, we predict the classification via an updated model.
-Discussion.
+Discussion
 In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally.
@@
 It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels.
-Our experiments in Sec.
-4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach.
+Our experiments in subsection 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach.
 Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.
 4 Experimental Results
-Datasets.
-The experiments here were conducted on three benchmark datasets: CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C [15], each containing a variety of common image corruptions.
-CIFAR-10-C includes 18 corruption types, while CIFAR-100-C features 19 types, covering weather, noise, and digital distortions.
-Both CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes of 32$\times$32 color images, respectively, similar to CIFAR-10 and CIFAR-100 [24].
-Tiny-ImageNet-C, derived from the Tiny-ImageNet dataset, contains 15 corruption types and 200 classes of 64$\times$64 color images.
-In addition, we also considered a neuromorphic dataset, DVS Gesture-C [18], and the related experiments are provided in the Supplementary Material Sec.
-7.
-Experimental Setup.
-All experiments were conducted under the highest corruption level (level=5).
-Following the experimental setup of MEMO [43], we applied AugMix [16] augmentation with a batch size of 32 to each test sample to generate a diverse set of augmented samples.
-The backbone model for most experiments is an SNN-based VGG architecture with Batch-Normalization Through Time (BNTT) [22], featuring BN layers at each time step, with 25 inference steps.
-To demonstrate the general applicability of the proposed SPACE method, additional experiments were conducted using an SNN-based ResNet-11 model [26] on the CIFAR-10-C dataset, with inference steps set to 30.
-Baseline TTA Methods.
-We compared our approach with models pre-trained on clean datasets without any test-time adaptation (referred to as w.o.
-TTA).
-Moreover, since few TTA methods are specifically designed for SNNs, we selected two most representative traditional TTA methods, MEMO [43] and SITA [21], as our baselines.
+Datasets and Models
+The experiments were conducted on six benchmark datasets for out-of-distribution generalization: CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], and ImageNet-A [18].
+The first three datasets include 15 corruption types across four categories and five severity levels, while the three ImageNet-based datasets capture real-world domain shifts, such as distributional discrepancies and semantic variations.
+Additionally, we evaluated on the neuromorphic dataset DVS Gesture-C [20], which introduces six corruption types.
+Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[24], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[27]; 3) Spike-driven Transformer V3 [44], containing 19M parameters; and 4) SNN-ConvLSTM.
+All models were pre-trained on their respective source domains to ensure fair evaluation on datasets with domain shifts.
+Compared Methods
+We compared our approach with models pre-trained on clean datasets and evaluated without any test-time adaptation (referred to as No Adapt).
+Additionally, since there are few TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [47] and SITA [23], as baselines.
+These methods represent the current state-of-the-art for source-free and single-instance TTA, respectively.
+MEMO enforces output consistency among augmented samples, making it naturally compatible with SNNs by directly adjusting their final outputs.
+SITA, designed for models with BN layers, was adapted to SNN-VGG with BNTT by updating BN layer parameters during TTA.
+Implementation details
+To maximize the effectiveness of our proposed method, all experiments conducted on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C , and DVS Gesture-C were performed with the highest severity level (level=5).
+In line with our setting and practical scenarios, the test batch size for all experiments was fixed to 1, and the model was updated only once per test sample.
+Following MEMO’s experimental setup, we applied AugMix [16] augmentation to generate diverse augmented samples, using a batch size of 32 for CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C, and 64 for the ImageNet series.
+Please refer to Appendix for more details.
 4.1 Main Results
-Shifted Domain w.o. TTA SITA MEMO SPACE
-noise gaussian 72.38% 73.06% 77.73% 77.98%
-shot 74.70% 74.15% 79.50% 79.34%
-speckle 71.75% 71.67% 77.28% 77.38%
-impulse 58.57% 58.41% 65.74% 69.41%
-blur defocus 63.05% 62.94% 65.61% 71.59%
-gaussian 57.03% 57.36% 59.93% 68.31%
-motion 64.44% 64.36% 66.24% 72.14%
-zoom 71.33% 70.72% 72.61% 74.67%
-weather snow 76.32% 76.67% 77.38% 78.43%
-fog 43.57% 43.40% 45.47% 52.80%
-frost 75.72% 75.24% 78.64% 79.59%
-digital brightness 82.44% 82.51% 83.05% 83.22%
-contrast 22.54% 22.02% 23.83% 23.85%
-elastic_transform 75.01% 74.74% 76.24% 75.49%
-pixelate 70.25% 69.62% 73.25% 76.24%
-jpeg_compression 84.28% 84.16% 86.05% 82.88%
-spatter 78.42% 78.61% 79.75% 77.52%
-saturate 80.85% 80.97% 82.03% 82.54%
-Average 67.93% 67.81% 70.53% 72.41%
-Table 1: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
-In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO [43] and SITA [21], which are current state-of-the-art methods for source-free and single-instance TTA.
-While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation.
-Similarly, SITA, despite not using backpropagation, fails to capture the intrinsic temporal dynamics of SNNs, rendering it ineffective.
-To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using the VGG architecture, and on CIFAR-10-C with the ResNet-11 architecture to demonstrate the broad applicability of our proposed method.
-4.1.1 Performance Comparison on VGG Model
-Shifted Domain w.o. TTA SITA MEMO SPACE
-noise gaussian 42.51% 43.01% 43.46% 44.71%
-shot 45.40% 44.83% 45.41% 46.73%
-speckle 43.35% 43.14% 43.53% 44.84%
-impulse 25.50% 25.48% 26.07% 27.99%
-blur defocus 42.15% 42.23% 42.47% 43.35%
-gaussian 37.76% 37.88% 38.29% 39.19%
-motion 41.87% 42.11% 42.55% 43.46%
-zoom 46.77% 46.93% 47.21% 48.42%
-glass 42.84% 43.29% 43.90% 44.47%
-weather snow 44.88% 44.49% 44.74% 45.70%
-fog 16.51% 16.65% 16.76% 17.62%
-frost 44.12% 43.84% 44.34% 44.98%
-digital brightness 49.36% 49.53% 49.95% 50.41%
-contrast 5.42% 5.56% 5.62% 5.69%
-elastic_transform 51.33% 50.99% 51.24% 51.45%
-pixelate 52.35% 52.48% 52.82% 53.81%
-jpeg_compression 58.47% 58.71% 59.29% 58.96%
-spatter 47.76% 47.82% 48.63% 48.95%
-saturate 46.91% 47.17% 47.17% 48.22%
-Average 41.33% 41.38% 41.76% 42.58%
-Table 2: Performance comparison of different methods on CIFAR-100-C dataset at level 5 corruption, using the VGG11 architecture.
-Shifted Domain w.o. TTA SITA MEMO SPACE
-noise gaussian 12.43% 12.71% 13.64% 16.71%
-shot 15.20% 15.17% 16.45% 19.44%
-impulse 8.74% 8.87% 9.51% 11.72%
-defocus 7.58% 7.56% 7.51% 9.62%
-blur motion 14.48% 14.57% 14.38% 16.57%
-zoom 13.59% 13.94% 13.92% 15.61%
-glass 6.50% 6.57% 6.60% 7.31%
-weather snow 15.58% 15.72% 15.97% 18.66%
-fog 5.27% 5.32% 4.91% 6.01%
-frost 18.06% 18.05% 17.83% 21.02%
-digital brightness 15.01% 14.82% 14.74% 17.53%
-contrast 1.32% 1.38% 1.44% 1.51%
-elastic_transform 20.53% 21.45% 21.00% 21.96%
-pixelate 31.62% 32.20% 31.39% 31.89%
-jpeg_compression 31.15% 31.51% 30.82% 31.63%
-Average 14.47% 14.66% 14.67% 16.48%
-Table 3: Performance comparison of different methods on Tiny-ImageNet-C dataset at level 5 corruption, using the VGG11 architecture.
-Shifted Domain w.o. TTA MEMO SPACE
-noise gaussian 72.88% 74.21% 75.90%
-shot 74.49% 75.50% 77.81%
-speckle 73.12% 74.32% 76.75%
-impulse 49.59% 51.35% 56.85%
-blur defocus 67.77% 68.11% 68.66%
-gaussian 63.57% 64.11% 64.15%
-motion 65.14% 65.37% 66.02%
-zoom 73.06% 73.96% 73.82%
-weather snow 78.11% 78.32% 78.74%
-fog 42.84% 43.47% 45.47%
-frost 73.12% 73.39% 74.04%
-digital brightness 81.71% 81.82% 81.28%
-contrast 12.72% 13.63% 14.94%
-elastic_transform 74.88% 75.54% 76.79%
-pixelate 77.08% 77.54% 80.32%
-jpeg_compression 82.47% 83.24% 83.96%
-spatter 78.75% 79.46% 80.58%
-saturate 79.40% 79.40% 78.70%
-Average 67.82% 68.49% 69.71%
-Table 4: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the ResNet11 architecture.
-The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Tab.
-1, Tab.
-2, and Tab.
-3, demonstrate the effectiveness of SPACE.
-SPACE outperforms w.o.
-TTA, MEMO, and SITA in average accuracy across all datasets.
-On CIFAR-10-C, SPACE achieves 72.41%, improving by 4.48% over w.o.
-TTA (67.93%), 4.60% over SITA (67.81%), and 1.88% over MEMO (70.53%).
-On CIFAR-100-C, SPACE reaches 42.58%, surpassing w.o.
-TTA (41.33%), SITA (41.38%), and MEMO (41.76%).
-On Tiny-ImageNet-C, SPACE achieves 16.48%, compared to 14.47% for w.o.
-TTA, 14.66% for SITA, and 14.67% for MEMO.
-Notably, SPACE consistently improves performance on Tiny-ImageNet-C across all domains, showing its robustness on complex datasets.
-Across individual corruption types, SPACE consistently outperforms w.o.
-TTA, MEMO, and SITA.
-For instance, on CIFAR-10-C, SPACE shows notable improvements, such as in impulse noise (69.41% for SPACE vs.
-58.57% for w.o.
-TTA, 58.41% for SITA, and 65.74% for MEMO) and fog (52.80% for SPACE vs.
-43.57% for w.o.
-TTA, 43.40% for SITA, and 45.47% for MEMO).
-On CIFAR-100-C, SPACE similarly excels, including in motion blur (43.46% for SPACE vs.
-41.87% for w.o.
-TTA, 42.11% for SITA, and 42.55% for MEMO).
-A particularly striking result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type.
-For example, in gaussian noise and impulse noise, SPACE leads by a significant margin.
-Even in simpler corruptions like brightness and pixelate, SPACE outperforms the w.o.
-TTA, demonstrating its robustness in complex datasets like Tiny-ImageNet-C.
-Although SPACE generally outperforms MEMO and SITA, MEMO does slightly better in a few cases, such as jpeg_compression on CIFAR-10-C (86.05% for MEMO vs.
-82.88% for SPACE) and on CIFAR-100-C (59.29% for MEMO vs.
+Table 1: Performance comparison on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C with CNN-based SNN models regarding Accuracy (%). The bold number indicates the best result.
+Method Noise Blur Weather Digital Average
+Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+CIFAR-10-C, SNN-VGG9
+No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
+SITA 73.06 74.15 58.41 62.94 64.21 64.36 70.72 76.67 43.40 75.24 82.51 22.02 74.74 69.62 84.16 66.41
+MEMO 77.73 79.50 65.74 65.61 67.50 66.24 72.61 77.38 45.47 78.64 83.05 23.83 76.45 73.25 85.05 69.20
+SPACE (ours) 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
+CIFAR-10-C, SNN-ResNet11
+No Adapt 72.88 74.49 49.59 67.77 63.53 65.14 73.06 78.11 42.84 73.12 81.71 12.72 74.88 77.08 82.47 65.96
+MEMO 74.21 75.50 51.35 68.11 64.65 65.37 73.96 78.32 43.47 73.39 81.82 13.63 75.54 77.54 83.24 66.67
+SPACE (ours) 75.90 77.81 56.85 68.66 67.64 66.02 73.82 78.74 45.47 74.04 81.28 14.94 76.79 80.32 83.96 68.15
+CIFAR-100-C, SNN-VGG11
+No Adapt 42.51 45.40 25.50 42.15 42.84 41.87 46.77 44.88 16.51 44.12 49.36 5.42 51.33 52.35 58.47 40.63
+SITA 43.01 44.83 25.48 42.23 43.29 42.11 46.93 44.49 16.65 43.84 49.53 5.56 50.99 52.48 58.71 40.68
+MEMO 43.46 45.41 26.07 42.47 43.90 42.55 47.21 44.74 16.76 44.34 49.95 5.62 51.24 52.82 59.29 41.06
+SPACE (ours) 44.71 46.73 27.99 43.35 44.47 43.46 48.42 45.70 17.62 44.98 50.41 5.69 51.45 53.81 58.96 41.58
+Tiny-ImageNet-C, SNN-VGG11
+No Adapt 12.43 15.20 8.74 7.58 6.50 14.48 13.59 15.58 5.27 18.06 15.01 1.32 20.53 31.62 31.15 14.47
+SITA 12.71 15.17 8.87 7.56 6.57 14.57 13.94 15.72 5.32 18.05 14.82 1.38 21.45 32.20 31.51 14.66
+MEMO 13.64 16.45 9.51 7.51 6.60 14.38 13.92 15.97 4.91 17.83 14.74 1.44 21.00 31.39 30.82 14.67
+SPACE (ours) 16.71 19.44 11.72 9.62 7.31 16.57 15.61 18.66 6.01 21.02 17.53 1.51 21.96 31.89 31.63 16.48
+In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO and SITA.
+To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using SNN-VGG, on CIFAR-10-C with SNN-ResNet11, on ImageNet series with Spike-driven Transformer V3, and on DVS Gesture-C with SNN-ConvLSTM, to demonstrate the broad applicability of our proposed method.
+Performance Comparison on CNN-based SNN Models
+Figure 3: Performance comparison on ImageNet-V2/R/A with Spike-driven Transformer V3 regarding Accuracy (%). The bold number indicates the best result.
+Method Accuracy (%)
+V2 R A Avg.
+No Adapt 67.64 41.94 15.12 41.57
+MEMO 67.57 42.15 14.97 41.56
+SPACE (ours) 68.82 45.07 16.22 43.37
+The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Table 1, highlight the effectiveness of SPACE.
+Key observations include the following: 1) Superior Performance Across Methods: SPACE consistently outperforms No Adapt, SITA, and MEMO in average accuracy across all datasets under both VGG and ResNet architectures.
+2) Broad Applicability: SPACE demonstrates effectiveness across different CNN structures, showcasing its generalizability.
+3) Robustness on Complex Datasets: A particularly notable result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type.
+This highlights its robustness in handling complex datasets like Tiny-ImageNet-C.
+4) Limitations of SITA on SNNs: The SNN-ResNet does not include BN layers, rendering SITA inapplicable.
+Furthermore, SITA performs worse than No Adapt, indicating its ineffectiveness for SNNs.
+This may be attributed to the spike-based transmission in SNNs: directly adjusting BN parameters without gradient guidance can disrupt spike information, leading to significant performance degradation.
+5) Comparison with MEMO: While SPACE generally outperforms MEMO and SITA, MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs.
+82.88% for SPACE) and CIFAR-100-C (59.29% for MEMO vs.
 58.96% for SPACE).
-These minor differences suggest MEMO has an edge with specific corruptions, but they do not diminish SPACE’s overall superiority.
-4.1.2 Performance Comparison on ResNet Model
-We evaluated the SPACE method on the CIFAR-10-C dataset using the ResNet11 architecture, which lacks BN layers, making SITA inapplicable.
-The results, shown in Tab.
-4, demonstrate SPACE’s effectiveness across different model structures.
-SPACE achieves an average accuracy of 69.71%, improving by 1.22% over MEMO (68.49%) and 1.89% over w.o.
-TTA (67.82%).
-In terms of individual corruption types, SPACE outperforms MEMO and w.o.
-TTA in most cases.
-Notable improvements include challenging corruptions such as impulse noise (56.85% vs.
-49.59% for w.o.
-TTA and 51.35% for MEMO) and contrast (14.94% vs.
-12.72% for w.o.
-TTA and 13.63% for MEMO).
-SPACE also achieves 80.32% on pixelate, outperforming MEMO (77.54%) and w.o.
-TTA (77.08%).
-These trends align with those observed in the VGG-based experiments, reaffirming SPACE’s consistency and generalizability.
-While MEMO marginally outperforms SPACE on zoom (73.96% vs.
-73.82%) and brightness (81.82% vs.
-81.28%), these differences are minimal and do not overshadow SPACE’s overall advantages.
-The results further confirm that SPACE is not dependent on a specific architecture and can adapt effectively to various backbone models.
-4.2 Ablation Study
-Figure 3: Local Feature Maps Based SPACE Outperforms Global Feature Maps Based SPACE.
-Local vs.
-Global Feature Map Similarity.
+These minor differences suggest MEMO may have an advantage under certain corruption types, but they do not detract from SPACE’s overall superiority.
+Performance Comparison on Spike-driven Transformer V3
+To explore the real-world impact of SPACE, we conducted tests on diverse datasets, including ImageNet-V2, ImageNet-R, and ImageNet-A.
+Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-Event Transformer V3.
+As shown in Figure 3, SPACE consistently outperforms MEMO, achieving a 1.80% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
+Performance Comparison on ConvLSTM-based SNN Model
+SNNs are inherently well-suited for processing event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency.
+The DVS-Gesture dataset [1] provides an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures.
+To further explore potential network architectures, we incorporated ConvLSTM.
+As shown in Table 2, our approach achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
+Table 2: Performance comparison on DVS Gesture-C with ConvLSTM-based SNN model regarding Accuracy (%). The bold number indicates the best result.
+Method Accuracy (%)
+DropPixel DropEvent RefractoryPeriod TimeJitter SpatialJitter UniformNoise Avg.
+No Adapt 53.03 54.55 36.74 31.06 92.05 91.67 59.85
+MEMO 56.06 55.30 36.74 30.68 90.91 91.67 60.23
+SPACE (ours) 57.20 57.95 52.27 43.18 92.42 92.42 65.91
+4.2 Ablation Studies
+(a) Local feature maps based SPACE outperforms global feature maps based SPACE.
+(b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
+Figure 4: Results of ablation study.
+Local vs. Global Feature Map Similarity
 We compared the performance of local versus global feature map similarity for our proposed method.
-The results in Fig.
-3 clearly illustrate the advantage of using similarity of local feature maps over global feature maps.
+The results in 4(a) clearly illustrate the advantage of using similarity of local feature maps over global feature maps.
 While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency.
-As shown in Fig.
-3, the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets.
+As shown in 4(a), the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets.
 This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation.
 These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
-Figure 4: Distribution of similarity values between spike counts based feature maps before and after SPACE adaptation.
-Effect of SPACE on Feature Map Similarity.
-To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Eq.
-4 $\sim$ Eq.
-6) across the CIFAR-10-C dataset.
-As shown in Fig.
-4, we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Gaussian Blur and Impulse Noise domains.
+Effect of SPACE on Feature Map Similarity
+To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset.
+As shown in 4(b), we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain.
 The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process.
-Specifically, the “After SPACE Adaptation” distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples.
+Specifically, the "After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples.
 This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
-5 Conclusion
+5 Conclusion and Limitation
 In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs.
-SPACE leverages SNNs’ unique spike dynamics to optimize spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods, which often ignore the temporal and sparse nature of SNN computation.
-Extensive experiments on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C show SPACE achieves consistent performance improvements under severe corruptions, while generalizing effectively to various architectures like VGG9 and ResNet11.
-These results demonstrate SPACE’s robustness and adaptability to domain shifts, paving the way for its potential use in neuromorphic hardware and energy-efficient AI systems.
-Supplementary Material
-6 Kernel-Based Consistency Regularization
+By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations.
+Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures, demonstrating robustness to domain shifts.
+However, SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios.
+Future work will explore alternative optimization strategies to expand its practicality for neuromorphic hardware and energy-efficient AI.
+Appendix A Kernel-Based Consistency Regularization
 To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function.
-The channel-wise feature probability distribution $\mathbf{P}_{c}$ obtained from Eq.
-4 can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
+The channel-wise feature probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
 $k(\mathbf{z},\mathbf{z}^{\prime})=\text{exp}(-\frac{\|\mathbf{z}-\mathbf{z}^{\prime}\|^{2}}{2\sigma^{2}}),$ (6)
@@
 $\mathcal{L}^{*}(\theta;\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (13)
-Shifted Domain w.o. TTA SPACE SPACE + Kernel Embedding
-noise gaussian 72.38% 77.98% 77.85%
-shot 74.70% 79.34% 79.34%
-speckle 71.75% 77.38% 77.33%
-impulse 58.57% 69.41% 69.25%
-blur defocus 63.05% 71.59% 71.42%
-gaussian 57.03% 68.31% 67.98%
-motion 64.44% 72.14% 72.21%
-zoom 71.33% 74.67% 74.61%
-weather snow 76.32% 78.43% 78.48%
-fog 43.57% 52.80% 52.46%
-frost 75.72% 79.59% 79.66%
-digital brightness 82.44% 83.22% 83.25%
-contrast 22.54% 23.85% 23.40%
-elastic_transform 75.01% 75.49% 75.42%
-pixelate 70.25% 76.24% 76.13%
-jpeg_compression 84.28% 82.88% 82.98%
-spatter 78.42% 77.52% 77.47%
-saturate 80.85% 82.54% 82.64%
-Average 67.93% 72.41% 72.41%
-Table 5: Performance comparison of SPACE without and with kernel embedding on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
-Shifted Domain w.o. TTA MEMO SPACE
-DropPixel 44.53% 43.94% 46.97%
-DropEvent 50.00% 51.52% 53.41%
-RefractoryPeriod 35.16% 34.85% 39.39%
-TimeJitter 42.19% 40.91% 42.42%
-SpatialJitter 83.98% 82.95% 85.23%
-UniformNoise 75.39% 74.24% 75.39%
-Average 55.21% 54.74% 57.14%
-Table 6: Performance comparison of different methods on DVS Gesture-C dataset at level 5 corruption.
-To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using the BNTT model with the VGG9 architecture, as shown in Tab.
-5.
+To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using SNN-VGG9 model, as shown in Table 3.
 The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical.
@@
 These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
-7 Performance on DVS Gesture-C Dataset
-SNNs are inherently suited for processing event-based neuromorphic data due to their temporal dynamics and energy efficiency.
-Static image datasets like CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C are commonly used to evaluate robustness to corruption and distribution shifts, but they lack the temporal and asynchronous characteristics of real-world event-based data.
-DVS-Gesture [1] allows us to evaluate the proposed method’s performance on neuromorphic datasets, emphasizing its compatibility with dynamic vision tasks and data with temporal structure.
-Following the state-of-the-art methodology [18], we evaluate the robustness improvements achieved by SPACE using DVS-Gesture-C, a corrupted variant of the standard DVS-Gesture dataset.
+Table 3: Performance comparison of SPACE without/with kernel embedding (k.e.) on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%). The bold number indicates the best result.
+Method Noise Blur Weather Digital Average
+Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
+SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
+SPACE + k.e. 77.85 79.34 69.25 71.42 67.64 72.21 74.61 78.48 52.46 79.66 83.25 23.40 75.42 76.13 82.98 70.94
+Appendix B Exploration of Alternative Alignment Objectives
+B.1 Two Alternative Alignment Objectives
+Average Membrane Potential (SPACE + a.m.p.)
+In the main paper, we use feature maps based on the spike counts over the entire time window as the alignment objective.
+In this section, we explore an alternative biologically inspired metric: the average membrane potential obtained from the LIF neurons in the final layer of the feature extractor.
+Specifically, for a given test point $\mathbf{x}_{i}$, the membrane potentials are captured in the feature map $\mathbf{U}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times D}$, where $T$, $C$, and $D$ represent the time steps, the number of channels, and the spatial dimensions, respectively.
+For LIF neurons that spike at a specific time step, we use the membrane potential after the LIF reset, which indirectly captures the spiking information.
+To obtain the primary objective feature map $\mathbf{F}(\mathbf{x}_{i})$, we compute the temporal average of $\mathbf{U}(\mathbf{x}_{i})$ along the first dimension.
+The resulting feature map is then used to calculate the loss following Equations 4 and Equation 5 in the main paper.
+Finally, the model is updated via SGD.
+Spikes through All Time Steps (SPACE + s.t.t.)
+Instead of summing the spike counts over the temporal dimension, we explicitly retain the temporal information, resulting in a feature map with an additional dimension that captures the spiking activity at each time step.
+This approach preserves the precise temporal dynamics of the spikes and leverages them directly for alignment.
+Specifically, we align the channel-wise spiking activity across all time steps between each pair of augmented samples.
+The update rules remain consistent with those described in the main paper.
+B.2 Experimental Results and Analysis
+Table 4: Performance comparison of SPACE with different objectives on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%).
+The bold number indicates the best result.
+Average membrane potential and spike through all time steps are referred as a.m.p.
+and s.t.t.
+respectively.
+Method Noise Blur Weather Digital Average
+Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
+SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
+SPACE + a.m.p. 74.87 76.05 61.64 64.90 65.79 65.88 72.22 77.66 45.66 76.64 82.81 21.57 75.53 70.91 84.74 67.79
+SPACE + s.t.t. 72.74 73.84 57.86 62.91 63.66 64.33 70.94 76.35 43.16 74.77 82.30 21.20 74.24 69.35 83.97 66.11
+We implement the two approaches above and evaluate on CIFAR-10-C dataset with SNN-VGG9 model. The results are shown in Table 4.
+Compared to No Adapt, SPACE + a.m.p.
+improves the generalization ability of the model.
+However, its performance is noticeably inferior to SPACE on the majority of corruptions.
+This may be due to the fact that the membrane potentials of the same sample under different augmentations tend to exhibit similar patterns, making them less effective in capturing the temporal distribution characteristics of SNNs.
+For example, in extreme cases where the membrane potential of a LIF neuron stays near the firing threshold throughout all time steps without crossing it, SPACE + a.m.p.
+fails to capture this information.
+In contrast, SPACE can reflect this behavior through spike counts, which more directly encode spiking activity.
+SPACE + s.t.t.
+fails to improve the generalization ability of the model.
+We hypothesize that this is due to the model overfitting during the alignment process.
+Specifically, the updates made to minimize the alignment loss may inadvertently lead to excessive specialization to the augmented samples, thereby hindering its ability to generalize effectively to unseen corruptions.
+Appendix C Additional Ablation Study
+Effect of Augmentation Quantity per Test Point
+In the main paper, we follow MEMO [47] to determine the number of augmentations $M$.
+To analyze the effect of $M$ on our method, we evaluated the average accuracy and the evaluation time per test point the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$.
+As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement.
+For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly.
+Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
+Table 5: Performance comparison of SPACE with different augmentation quantities on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%) and Evaluation Time (seconds) per test point.
+No Adapt Number of Augmentation
+2 4 8 16 32 64 128
+Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
+Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
+Appendix D Additional Experimental Information
+D.1 More Details on Datasets
+In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability.
+They are CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], ImageNet-A [18] and DVS Gesture-C [20].
+CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
+These three datasets are commonly used benchmarks for evaluating the robustness of models under various types of input corruptions.
+They are derived from the original CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets by applying 15 corruption types (e.g., noise, blur, weather, and digital distortions) at 5 severity levels to the original CIFAR-10, CIFAR-100 [26], and Tiny-ImageNet [6] datasets, respectively.
+CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes, each with 50,000 training images and 10,000 testing images.
+Tiny-ImageNet-C, derived from Tiny-ImageNet, includes 200 classes with 100,000 training images and 10,000 validation images.
+ImageNet V2/R/A
+ImageNet-V2, ImageNet-R, and ImageNet-A are three datasets designed to evaluate model robustness under different real-world challenges.
+ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [6], serving as a test of distribution shift.
+ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes.
+ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses.
+Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
+DVS Gesture-C
+DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset.
 This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise.
 These corruptions, implemented via the Tonic API [29], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
-To ensure a comprehensive and stringent assessment, we consistently employ the highest severity level across all experimental evaluations.
-The pre-trained model used in this study follows a 2CONV-2FC architecture, which was selected for its balance between computational efficiency and representational capacity, ensuring a fair and reliable comparison.
-The experimental results demonstrate that the proposed SPACE method achieves superior performance on the DVS-Gesture-C dataset compared to both w.o.
-TTA and the MEMO method.
-As shown in Tab.
-6, SPACE consistently outperforms w.o.
-TTA across all six corruption types, with the largest improvements observed for RefractoryPeriod (+4.23%) and DropEvent (+3.41%).
-Compared to MEMO, SPACE also achieves higher accuracy in most scenarios, with notable gains for RefractoryPeriod (+4.54%) and DropPixel (+3.03%).
-On average, SPACE achieves an accuracy of 57.14%, significantly surpassing both w.o.
-TTA (55.21%) and MEMO (54.74%).
-While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation.
-This limitation is evident in the results, where MEMO struggles to handle certain corruption types effectively, such as RefractoryPeriod and TimeJitter.
-In contrast, SPACE addresses these challenges by fully leveraging the temporal dynamics and event-based nature of SNNs, resulting in more robust and reliable performance across diverse corruption types.
-These results highlight the robustness and effectiveness of SPACE in improving model performance under challenging conditions, making it a highly suitable method for event-based data with various noise and corruption types.
+D.2 More Details on Pretrained Models
+SNN-VGG with BNTT
+We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [24] as our main backbone model.
+BNTT decouples the parameters along the time axis within each layer, effectively capturing the temporal dynamics of spikes.
+The temporally evolving learnable parameters in BNTT enable neurons to regulate their spike rates across different time steps, facilitating low-latency and low-energy training from scratch.
+The pre-trained model achieves accuracies of 90.61%, 69.50% and 58.36% on the original CIFAR-10, CIFAR-100 and Tiny-ImageNet test sets respectively.
+SNN-ResNet
+To explore other CNN architecture, we use SNN-ResNet11 [27], which trains SNNs directly via surrogate gradient.
+The pre-trained model used in our experiments can achieve 90.95% accuracy on the original CIFAR-10 test sets.
+Spike-driven Transformer V3
+To evaluate the effectiveness of our work on more complex networks and datasets, we utilized a 19M parameter model in Spike-driven Transformer V3 [44], which achieves 79.60% accuracy on the ImageNet-1K validation set.
+This model has a patch size of 14 $\times$ 14, with an embedding dimension of 360 output by the final feature extractor.
+Instead of using the class token, the entire embedding feature map is employed as the input to the classifier.
+This design naturally aligns with our method, where the channel dimension $C$ corresponds to the embedding dimension, and the spatial dimension $D$ matches the patch size.
+SNN-ConvLSTM
+DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures.
+To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64, respectively.
+This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
+D.3 More Details on Implementation
+To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input.
+In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings.
+For each set of experiments, we employ the SGD optimizer and evaluate learning rates ranging from 0.001 to 0.8, selecting the optimal value based on performance.
+We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.

### v3

@@
 We evaluate SPACE on multiple datasets.
-Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs (such as VGG and ResNet), Transformer models, and ConvLSTM architectures.
-Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
+Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs, Transformer, and ConvLSTM architectures.
+Experimental results show that SPACE outperforms state-of-the-art ANN methods while maintaining lower computational cost, highlighting its effectiveness and robustness for SNNs in real-world settings.
+The code will be available at https://github.com/ethanxyluo/SPACE.
 1 Introduction
-Recent advancements in neuroscience-inspired computing have placed Spiking Neural Networks (SNNs) at the forefront of research as a biologically plausible alternative to traditional Artificial Neural Networks (ANNs).
-While ANNs have achieved remarkable success in various domains, their reliance on dense, black-box architectures often limits their interpretability and energy efficiency.
-In contrast, SNNs emulate the sparse, event-driven dynamics of biological neurons, providing several advantages in terms of computational efficiency, temporal processing, and explainability [4, 12, 37].
-However, the unique features that make SNNs advantageous—such as their reliance on temporal information and sparse spiking—also pose significant challenges in dynamic real-world environments.
-SNNs encode information dynamically across time, making them highly sensitive to changes in input data distribution [2].
-This sensitivity can result in drastic changes to spiking behavior across the network, ultimately degrading performance when faced with distribution shifts.
-Understanding and enhancing the robustness of SNNs under such conditions is therefore a critical research direction.
-In real-world applications, it is common that the distribution of the training data (a.k.a.
-source domain) and the test data (a.k.a.
-target domain) are different due to variations in lighting, background, sensor noise, or environmental conditions.
-This phenomenon, known as domain shift, can significantly degrade the performance of machine learning models [15, 25].
-SNNs are particularly vulnerable to domain shifts, as even minor changes in input data can disrupt the temporal dynamics and spiking behavior in intermediate layers, leading to poor generalization.
-Traditional test-time adaptation (TTA) solutions such as retraining or fine-tuning on target domain data with help of source data related proxy task [41], statistics of source data [33], or a batch of test data [42], are often impractical in real-world scenarios due to the unavailability of source data or other test samples, particularly in privacy-sensitive or resource-constrained environments.
-Source-free and single-instance TTA methods [9, 23, 47] offer a promising alternative by enabling the model to adapt to the target domain during inference without access to source data, which is typically defined (and also in this work) as one iteration of training adaptation with one test sample.
-Among existing source-free and single-instance TTA methods, MEMO [47] and SITA [23] are two representative approaches, but both face fundamental limitations when applied to SNNs.
-MEMO optimizes the entropy of the averaged prediction across multiple augmented views of a test sample, enforcing output probability consistency.
-However, this strategy may not be optimal for SNNs.
-In SNNs, output probability distributions reflect only high-level semantic information and fail to capture the fine-grained spiking dynamics that encode essential features which prevents MEMO from effectively mitigating domain shifts within SNNs’ internal representations.
-On the other hand, SITA adapts batch normalization (BN) statistics by shifting them toward those computed from test-time augmentations, but this approach is also unsuitable for SNNs.
-In SNNs, many models lack BN layers [27, 46], which limits the applicability of BN-based adaptation methods such as SITA in SNNs.
-Additionally, due to the sparsity of SNN activations, modifying BN statistics does not significantly alter neuronal spiking patterns, further reducing its ability to facilitate effective adaptation.
-In this work, we propose a novel method names SPike-Aware Consistency Enhancement (SPACE), which is the first source-free and single-instance TTA approach specifically designed for SNNs.
+Recent advancements in neuroscience-inspired computing have placed Spiking Neural Networks (SNNs) at the forefront as a biologically plausible alternative to traditional Artificial Neural Networks (ANNs).
+While ANNs have achieved remarkable success across domains, their reliance on dense, black-box architectures often limits interpretability and energy efficiency.
+In contrast, SNNs emulate the sparse, event-driven dynamics of biological neurons, offering advantages in computational efficiency, temporal processing, and explainability [13, 4, 44].
+However, the very features that make SNNs appealing—temporal coding and sparse spiking—also introduce fragility in dynamic real-world environments.
+Because SNNs encode information over time, they can be highly sensitive to shifts in the input distribution [2], which may induce substantial changes in layer-wise spiking activity and degrade downstream performance.
+At the same time, the historical SNN–ANN accuracy gap on standard vision benchmarks has narrowed considerably [58, 52], foregrounding generalization and robustness as primary objectives for practical SNN deployment [10, 51].
+In application domains central to SNNs—event-driven sensing and low-power edge inference—the ability to remain robust to distribution shift is thus a deployment prerequisite rather than a post-hoc enhancement.
+In real-world deployments, it is common for the training (a.k.a.
+source) and test (a.k.a.
+target) distributions to differ owing to changes in lighting, background, sensor noise, or environmental conditions.
+Such domain shifts are known to impair model performance [18, 30].
+Prior studies have specifically noted robustness challenges of SNNs under distribution shift [38, 27, 23], but remain hard to deploy online at scale.
+Mei et al.
+[38] perform gradual, unsupervised adaptation via teacher–student self-training across intermediate domains, which requires iterative weight updates, pseudo-label calibration, and multiple passes during deployment—assumptions that clash with single-sample, low-latency edge constraints.
+Karilanova et al.
+[27] address only temporal-resolution shifts via parameter remapping that presumes a known source–target sampling ratio and synchronized binning, offering no per-instance online adaptation and limited coverage of broader appearance/sensor shifts.
+To address the limitations, we adopt SNN-specific test-time adaptation (TTA) to meet practical deployment constraints.
+Existing TTA approaches often rely on source data or batched target samples—e.g., using source-related proxy tasks [48], source statistics [40], or optimization over a batch of test inputs [49, 54, 34].
+Such assumptions are frequently impractical due to privacy, regulatory, and on-device resource constraints.
+We therefore focus on the source-free, single-instance regime, where the model adapts online using only the current test sample and without access to source data or target batches.
+Note that several influential ANN-oriented TTA methods fall outside this regime.
+For example, SHOT [37] requires mini-batches of target data and multi-pass self-training in a static feature space while freezing a standalone classifier head; in SNNs, classification is tightly coupled to time-evolving spiking dynamics, leaving no stationary head to freeze and rendering distance-based pseudo-labeling insensitive to temporal structure.
+Similarly, TAST [25] relies on class-prior regularization and batch-level target statistics and is typically optimized over target mini-batches—assumptions misaligned with sparse, discrete spike trains and incompatible with the single-instance constraint.
+Within the truly source-free, single-instance family, MEMO [56] and SITA [28] are representative but face fundamental limitations for SNNs.
+MEMO enforces output-consistency by minimizing the entropy of the averaged prediction over augmented views; however, output probabilities capture only coarse semantics and provide weak control over the fine-grained temporal spiking dynamics that govern internal representations.
+SITA adapts batch-normalization (BN) statistics using test-time augmentations, yet many SNN architectures omit BN layers [55, 32]; even when present, sparse activations limit the effect of BN-moment shifts on spike generation.
+These gaps motivate an SNN-specific TTA framework that explicitly engages temporal spiking dynamics, avoids reliance on BN, and respects the source-free, single-instance constraints to improve robustness under domain shift.
+In this work, we propose a novel method named SPike-Aware Consistency Enhancement (SPACE), which is the first source-free and single-instance TTA approach specifically designed for SNNs.
 SPACE performs adaptation using only a single test point without access to source data, making it particularly suitable for real-world scenarios where source data are unavailable or privacy-sensitive.
 By leveraging the inherent spike dynamics of SNNs, SPACE maximizes the consistency in spike-behavior-based local feature maps across augmented samples of the same input, ensuring robust and efficient adaptation in deep SNNs.
-Unlike existing TTA methods, which often overlook the unique characteristics of SNNs, SPACE directly exploits spike-based representations for adaptation.
+Unlike existing TTA methods, which often overlook the unique characteristics of SNNs and the demanding of single instance scenarios, SPACE directly exploits spike-based representations for adaptation.
 Our contributions can be summarized as follows
   1. 1.
-    To the best of our knowledge, we are the first TTA method tailored for SNNs, which operates effectively using only a single test sample.
-This addresses the unique challenges of adapting SNNs in scenarios where source data are unavailable, while maintaining efficiency and scalability.
+    To the best of our knowledge, we are the first TTA method tailored for SNNs, which operates using only a single test sample.
+This addresses the unique challenges of adapting SNNs in scenarios where source data are unavailable, while maintaining efficiency and accuracy.
   2. 2.
-    Our method introduces a consistency-driven optimization framework to enhance the similarity of spike-behavior-based local feature maps across augmented samples.
+    Our method introduces a consistency-driven optimization framework to enhance the similarity of spike-behavior-based local feature maps across augmented samples of the single point.
 By leveraging spike dynamics, this approach ensures robust adaptation in deep SNNs.
@@
     To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [31], Tiny-ImageNet, ImageNet [7], and the neuromorphic dataset DVS Gesture [1].
-Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [24, 40], SNN-ResNet11 [14, 27], Spike-driven Transformer V3 [44], and SNN-ConvLSTM [39].
+Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [47, 29], SNN-ResNet [17, 32], Spike-driven Transformer V3 [52], and SNN-ConvLSTM [46].
 The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@
 [41] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance.
-These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [27, 43] and spiking recurrent networks [45, 48], which have denoted promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
-A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations.
-Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics.
-According to our experimental results presented in section 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time.
-Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
+These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [50, 32], spiking recurrent networks [53, 57], and spiking transformer networks [58, 52], which have denoted promising results in diverse tasks like image classification, speech recognition, and event-driven sensor data analysis.
 2.2 Test-time Adaptation
@@
 [33] employs pseudo-labeling to adjust model predictions.
-Although these approaches eliminate the need for source data, they often assume batch-level adaptation, relying on test-time statistics computed over multiple samples.
-This assumption limits their applicability in real-world settings where only single test points are available at a time.
-MEMO [47] addresses this scenario, proposing minimizing the marginal output distribution in the increases of the single test point.
-The objective in MEMO encourages the model to produce consistent predictions across different augmentations, thus enforcing the invariances encoded in these augmentations, while also maintaining confidence in its predictions.
-Another TTA method, SITA [23], adapts batch normalization statistics at inference by shifting them toward those computed from a batch of augmented versions of the single test instance.
-MEMO has shown that enforcing consistency among augmented samples improves TTA performance.
-However, in deep SNNs, where information is transmitted through discrete spikes, backpropagation tied to output objectives suffers from temporal information loss, making effective TTA challenging.
-Moreover, MEMO fails to capture the domain shift-induced variations in the internal behavior of the SNN, such as changes in neuron activations and spike patterns, which are critical for robust adaptation in SNNs.
-Although SITA avoids backpropagation, it also struggles to accurately model the intrinsic temporal dynamics of SNNs, limiting its ability to align augmented samples effectively.
-To overcome these issues, we enforce consistency by directly aligning the spike patterns and behaviors of augmented samples within the SNN framework.
-Unlike BN-dependent methods like SITA, our approach works regardless of the presence of BN layers, ensuring broader compatibility with various SNN architectures.
+Although these methods avoid source data, many assume batch-level adaptation, relying on test-time statistics over multiple samples—an impractical requirement when only a single test point is available.
+MEMO [56] addresses the single-instance setting by minimizing entropy and enforcing prediction consistency across augmentations, thereby promoting augmentation-invariant semantics but offering limited control over temporal spiking dynamics under shift.
+SITA [28] adapts batch-normalization statistics using augmented views of the same instance; however, many SNNs lack BN [55, 32], and sparse activations diminish the effectiveness of moment updates, especially under severe corruptions.
+To overcome these limitations, we align spike patterns and temporal features across augmentations without relying on BN, yielding an SNN-specific TTA mechanism compatible with diverse architectures and truly single-sample operation, while offering a more stable and informative alignment signal for robustness.
+2.3 Generalization of SNNs
+Recent state-of-the-art SNNs [58, 52] achieve performance on clean data comparable to strong ANNs, shifting the research focus from in-distribution accuracy to generalization and robustness under real-world shifts.
+Several works have begun to examine these issues in SNNs [10, 15, 51].
+Guo et al.
+[15] study Source-Free Domain Adaptation (SFDA) in a batch/epoch setting, where the model is updated after processing target-domain batches and typically assumes access to a pre-collected target dataset.
+In contrast, our work tackles TTA in a stricter, online regime: we adapt to each incoming sample individually and on the fly, without labels, source data, or target-set pre-collection.
+This per-sample update schedule reduces latency and memory footprint, making the approach directly applicable to streaming, autonomous deployments where SNNs are expected to excel.
 3 Proposed Method
 3.1 Preliminary
-SNNs emulate biological neurons by transmitting and encoding information through discrete spikes.
+SSNNs emulate biological neurons by transmitting and encoding information through discrete spikes.
 Among various neuron models, the Leaky Integrate-and-Fire (LIF) neuron [6] is widely adopted for its balance between biological realism and computational efficiency.
@@
 $u^{t}_{i}=(1-\frac{1}{\tau_{m}})u_{i}^{t-1}+\frac{1}{\tau_{m}}\sum_{j}w_{ij}o_{j}^{t}.$ (2)
+Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for connections between pre- and post-neurons.
+Figure 1 shows the LIF model, depicting membrane potential evolution and spike generation in response to input spikes [11].
 Figure 1: Illustration of spike and membrane potential dynamics in LIF neurons.
 When the membrane potential reaches the threshold, it is reset by subtracting the threshold value, triggering the neuron to fire a spike.
-Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons.
-Figure 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
 The spiking mechanism in SNNs introduces non-differentiability, making it difficult to apply standard gradient-based optimization methods for training or test-time adaptation.
@@
 $\frac{\partial o_{j}}{\partial U}\triangleq\begin{cases}1,&\text{if }U\geq U_{th},\\ 0,&\text{if }U<U_{th}.\end{cases}$ (3)
-While this is not an exact solution, it is valid because a reset occurs after a spike is generated when $U\geq U_{th}$.
-These foundational techniques for SNN training, including LIF dynamics and surrogate gradient approximations, form the basis for our proposed method, which focuses on enhancing TTA by ensuring consistency in spiking behavior across augmented samples.
+While this is not an exact analytical solution, it is nevertheless valid because a reset promptly occurs after a spike is generated when $U\geq U_{th}$.
+These SNN foundational techniques underpin our method to enhance TTA by ensuring consistent spiking behavior across augmented samples.
 3.2 SPike-Aware Consistency Enhancement
@@
 Algorithm Design Overview
-The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Figure 2.
+The proposed algorithm through spike-aware consistency enhancement (SPACE) for TTA is designed to improve test-time robustness of SNNs, as shown in Figure 2.
 We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample.
@@
 The overall method is presented in Algorithm 1, and the details are introduced below.
-The code will be made available after acceptance.
 Spike-Aware Feature Maps
@@
 It is worth noting that this is a reasonable assumption, supported by a recent advance [43] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware.
+Concretely, we decompose the SNN into a spike-based feature extractor and a classifier, writing $M_{\theta}(\mathbf{x})=C_{\theta_{C}}\!\big(E_{\theta_{E}}(\mathbf{x})\big)$ with $\theta=(\theta_{E},\theta_{C})$.
+Here, the extractor $E_{\theta_{E}}$ denotes the main body of the network (e.g., CNN/Transformer/ConvLSTM blocks), and the classifier $C_{\theta_{C}}$ consists of the last one or two readout layers.
 Aligned with previous work [28, 56], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective.
-In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples =$\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$.
+In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples $\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$.
 Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts.
-Different from MEMO [47], which adapts the model during test time via the conditional output distribution $p_{\theta}(y|\mathbf{x})$ of the model’s last layer, SPACE leverages the dynamic behavior of intermediate layers in the SNN that plays a crucial role in determining the temporal information and semantic representation of the final output.
-SNNs encode information through temporal dynamics and spike sparsity.
-These behavioral characteristics are particularly prominent in the intermediate layers and can be represented by feature maps $\mathbf{F}(\mathbf{x})$.
-Unlike ANNs, these characteristics in SNNs cannot be fully captured by the output probability distribution alone.
-Another advantage of intermediate feature maps is its highly sensitivity to changes in the input distribution due to the event-driven nature of SNNs.
-When the input data undergoes a distribution shift, the spiking patterns in the intermediate layers may change significantly, thereby impacting the final output.
-Optimizing the consistency of spiking behavior in intermediate layers can effectively mitigate feature variations caused by distribution shifts, thereby enhancing the robustness of models.
-To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window.
-As described in subsection 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$.
-The temporal spike patterns of the network are captured in the feature map $\mathbf{O}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps.
-$\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch.
-Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network, depending on whether the final feature map is pooled into a 1$\times$1 matrix.
-Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
-There are several reasons for choosing the total spike counts as the observation metric.
-First, SNNs typically operate over a time window ranging from 25 to thousands time steps [4, 13, 27, 34, 38], meaning that spike patterns across many time steps may be identical due to the repetitive nature of the spike dynamics.
-By comparing the cumulative spike counts instead of individual time-step patterns, we can effectively capture the temporal accumulation of spike activity while reducing redundancy.
-Second, the total spike counts reflects differences between augmented samples based on their overall spike dynamics, which is essential for our goal of aligning feature maps across augmentations.
-Finally, this representation significantly reduces the computational cost, as it condenses the temporal information into an aggregate value without requiring a step-by-step comparison of spike patterns.
+SPACE differs from MEMO [56] by operating not on the conditional output $p_{\theta}(y|\mathbf{x})$ but on the spike dynamics of intermediate layers that determine temporal coding and semantic representations.
+Specifically, we observe the feature map $\mathbf{O}(\cdot)\in\mathbb{R}^{T\times C\times H\times W}$ at the deepest layer of $E_{\theta_{E}}$ that still retains spatial support.
+Here $T$ denotes the number of time steps, $C$ represents the number of channels, $H$ and $W$ are the spatial dimensions of the local feature maps.
+If the last pre-classifier feature map has shape with $H\times W>1$, we use this layer; if global pooling or stride reduces it to $1\times 1$, we use the penultimate layer of $E_{\theta_{E}}$.
+This choice balances semantic abstraction and alignment granularity—using a $1\times 1$ map yields too few alignment targets and risks overfitting, whereas too shallow a layer may overemphasize low-level variations.
+To leverage the intermediate spike activity over the entire time window, let $\mathbf{O}(\mathbf{x}_{i})\in\{0,1\}$ denote binary spikes (1=fire, 0=silent, refers to subsection 3.1) at the chosen extractor layer for an augmented view $\mathbf{x}_{i}$.
+We aggregate spikes over time to form a spike-aware feature maps collection $\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
+We choose total spike counts for three main reasons: 1) SNNs often run for tens to thousands of steps [45, 16, 32, 41], making stepwise matching redundant and costly.
+2) Aggregation smooths the loss landscape: if spikes jitter in time but counts match, the loss remains unchanged, so gradients issue a clear directive—make this neuron more or less excitable for this input—producing stable, macro-level alignment of core feature activations across the augmented views.
+3) distribution shift primarily manifests as changes in spiking rates and spatial activation support in intermediate layers; $\mathbf{F}(\cdot)$ provides a direct, informative summary of these changes that output probabilities cannot capture.
 Algorithm 1 SPACE for Test-Time Adaptation in SNNs
-1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, model $\theta$, learning rate $\eta$
-2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the updated model
+1:Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, SNN model with extractor parameters $\theta_{E}$, and learning rate $\eta$
+2:Output: Prediction $y^{*}$ via updated model
 3:Step 1: Generate augmented batch
@@
 8:Step 2: Extract feature maps
-9:  Pass $\mathcal{B}$ through SNN, obtain =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
+9: Pass $\mathcal{B}$ through the SNN feature extractor, obtain $\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
 10:Step 3: Compute similarity
@@
 14:Step 4: Update model parameters
-15:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta$ via SGD with $\eta$
+15: Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta_{E}$ via SGD with learning rate $\eta$
 16:Step 5: Predict
-17:  $y^{*}=\arg\max_{y}p_{\theta^{*}}(y|\mathbf{x})$
-18:  Return: $y^{*}$
+[…]
+18:Return: Prediction result $y^{*}$
 Feature Maps Alignment
-The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data.
-When the feature maps of augmented samples are highly similar, the model becomes less sensitive to variations at test time, thereby improving its ability to generalize to new, unseen, and unlabeled data.
-To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$.
-Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ denotes the local feature vectors associated with channel $c$ of the augmented sample $\mathbf{x}_{i}$.
-Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$.
-This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively.
-To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
-$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
-The similarity measures the degree of overlap between the feature distributions, with higher values indicating greater consistency in spike-based representations across augmentations.
-Based on this similarity measure, we define the loss function for model adaptation given as
-$\mathcal{L}(\theta;\mathbf{x})\triangleq\sum_{1\leq j<i\leq M}\left(1-\bar{\mathcal{S}}(i,j|\mathbf{x})\right).$ (5)
-Then, we adapt the model using stochastic gradient descent (SGD) for a single iteration with the learning rate $\eta$.
-By minimizing this loss function, we enforce consistency across the feature maps of different augmentations.
-This reduces the variability induced by augmentations, ensuring that the model learns stable and robust feature representations.
-Furthermore, by promoting feature consistency, this approach serves as a regularization mechanism, preventing over-fitting to specific augmentations and encouraging the model to focus on more fundamental and generalizable features.
-While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure.
-Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) in Appendix).
-However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
-Finally, we predict the classification via an updated model.
-Discussion
-In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally.
-Since channels may have different active neurons, directly comparing the entire feature map fails to account for this variability.
-Instead, computing the similarity per channel and averaging the results focuses on the stability of feature representations within each channel.
-This isolates the contributions of active neurons, ensuring the similarity measure reflects meaningful changes rather than noise.
-It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels.
-Our experiments in subsection 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach.
-Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.
+Our goal is to make the model focus on invariant, task-relevant features rather than noise or minor transformations introduced by augmentations or distribution shift.
+If the feature maps of augmented views are highly similar, the model becomes less sensitive at test time and generalizes better to unseen, unlabeled samples.
+Concretely, given the spike-aware feature map $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ of an augmented view $\mathbf{x}_{i}$, we compute channel-wise local vectors $\mathbf{F}_{c}(\mathbf{x}_{i})\in\mathbb{R}^{D}$ and normalize them across spatial locations with a softmax to obtain a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})\in\Delta^{D-1}$.
+This normalization emphasizes salient regions and suppresses noise, yielding a stable target distribution.
+We then measure similarity between two augmented views $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ by the average channel-wise inner product
+$\bar{\mathcal{S}}(i,j\mid\mathbf{x})\;=\;\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\,\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
+which lies in $[0,1]$ and attains 1 if and only if the per-channel spatial distributions match exactly. Based on this, we adapt the extractor parameters by minimizing the loss function
+$\mathcal{L}(\theta_{E};\mathbf{x})\;\triangleq\;\sum_{1\leq j<i\leq M}\bigl(1-\bar{\mathcal{S}}(i,j\mid\mathbf{x})\bigr),$ (5)
+using a single SGD step with learning rate $\eta$.
+Minimizing $\mathcal{L}$ enforces consistency across augmentations, regularizing the model to learn stable, robust representations and preventing overfitting to any specific view.
+Finally, predictions are obtained with the adapted model.
+Theoretical Motivation
+Our objective instantiates consistency regularization at inference time: under semantics-preserving, bounded-severity augmentations (e.g., AugMix [19]), we enforce invariance in internal representations across transformations.
+Information-theoretically, aligning $\mathbf{F}$ (and thus $\mathbf{P}$) suppresses augmentation-specific variability while preserving identity-relevant content, approximating an information bottleneck that favors invariant features; geometrically, when perturbations push a sample off the learned manifold, adaptation pulls it toward the local manifold spanned by its augmented views, effectively denoising the representation.
+In SNNs, shifts often manifest as event-rate changes and timing jitter that nudge membrane potentials $U$ around the threshold $U_{th}$, making spike generation highly sensitive; by aligning time-resolved features across augmentations, we reduce the variance of $U$ trajectories and spike counts, enlarge the effective margin $|U-U_{th}|$, and stabilize and spike timing dynamics.
+Consequently, robustness arises from coupling semantic consistency with temporal stabilization within the SNN-TTA framework.
+We also explored richer feature relations via kernel embeddings and Maximum Mean Discrepancy (MMD) (Appendix B).
+In practice, the above linear-kernel objective already captures the essential distributional overlap with lower variance and cost, and stronger kernels did not bring noticeable gains while adding overhead.
+More discussion is presented in Appendix A.
 4 Experimental Results
+Table 1: Performance of different model architecture under domain shift.
+Values are Top-1 accuracy (%); Acc.
+Loss = Clean - Shifted.
+“Shifted” denotes evaluation on shifted variants of each dataset; exact configuration described in Appendix C.
+SNN and ANN models share backbone and scale and are trained with the same protocol.
+Architecture Dataset ANN SNN
+Clean Shifted Acc. Loss Clean Shifted Acc. Loss
+VGG9 CIFAR-10 90.53 77.25 13.28 90.61 67.86 22.74
+VGG11 CIFAR-100 70.86 48.32 22.54 69.50 41.35 28.14
+VGG11 Tiny-ImageNet 61.24 19.76 41.48 58.36 14.56 43.80
+Transformer ImageNet 85.16 75.49 9.67 79.60 67.64 11.96
 Datasets and Models
@@
 Additionally, we evaluated on the neuromorphic dataset DVS Gesture-C [23], which introduces six corruption types.
-Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[24], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[27]; 3) Spike-driven Transformer V3 [44], containing 19M parameters; and 4) SNN-ConvLSTM.
+Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[29], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[32]; 3) Spike-driven Transformer V3 (19M parameters) with 4 inference steps [52]; and 4) SNN-ConvLSTM with 32 inference steps.
 All models were pre-trained on their respective source domains to ensure fair evaluation on datasets with domain shifts.
 Compared Methods
-We compared our approach with models pre-trained on clean datasets and evaluated without any test-time adaptation (referred to as No Adapt).
-Additionally, since there are few TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [47] and SITA [23], as baselines.
+We compared our approach against models pre-trained on clean datasets and evaluated without any test-time adaptation (henceforth referred to as No Adapt), which serves as a lower-bound reference for robustness under distribution shift.
+Given the limited availability of TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [56] and SITA [28], as baselines.
 These methods represent the current state-of-the-art for source-free and single-instance TTA, respectively.
-MEMO enforces output consistency among augmented samples, making it naturally compatible with SNNs by directly adjusting their final outputs.
-SITA, designed for models with BN layers, was adapted to SNN-VGG with BNTT by updating BN layer parameters during TTA.
+MEMO enforces output consistency across augmented samples, making it inherently compatible with SNNs by directly adjusting their final outputs.
+SITA, originally designed for models with BN layers, was adapted for SNN-VGG with BNTT by updating the BN layer parameters during TTA.
+Additionally, we included RoTTA [54] and the recent DeYO [34] as baselines, despite their reliance on batch data.
+To ensure a fair comparison, we applied the same augmentation strategy as SPACE to construct a pseudo-batch for each sample.
 Implementation details
@@
 Following MEMO’s experimental setup, we applied AugMix [19] augmentation to generate diverse augmented samples, using a batch size of 32 for CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C, and 64 for the ImageNet series.
-Please refer to Appendix for more details.
+Please refer to Appendix C for more details.
+4.1 Evaluation of Robustness of SNNs under Domain Shift.
+To further substantiate the observation that SNNs are particularly vulnerable to domain shifts, we conducted a controlled comparison between SNNs and ANNs that share the same backbone family and model scale, trained under identical protocols, and evaluated on shifted variants of standard benchmarks (details in Appendix C).
+As summarized in Table 1, SNNs consistently incur larger accuracy drops than their ANN counterparts across CIFAR-10, CIFAR-100, Tiny-ImageNet, and ImageNet.
+These empirical results motivate the development of robust, SNN-specific TTA strategies tailored to spiking computation.
 4.2 Main Results
@@
 No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
+RoTTA 73.60 75.49 61.61 66.08 67.53 67.75 72.40 74.34 51.54 76.35 80.98 23.01 74.23 68.60 83.29 67.79
+DeYO 73.97 76.16 62.45 66.29 67.37 67.60 73.52 74.90 50.68 76.23 81.30 20.90 75.19 70.44 82.85 67.99
 SITA 73.06 74.15 58.41 62.94 64.21 64.36 70.72 76.67 43.40 75.24 82.51 22.02 74.74 69.62 84.16 66.41
@@
 No Adapt 72.88 74.49 49.59 67.77 63.53 65.14 73.06 78.11 42.84 73.12 81.71 12.72 74.88 77.08 82.47 65.96
+RoTTA 73.42 75.66 49.93 67.89 63.79 64.88 72.75 78.14 44.34 73.59 80.36 12.90 74.68 78.25 83.12 66.25
 MEMO 74.21 75.50 51.35 68.11 64.65 65.37 73.96 78.32 43.47 73.39 81.82 13.63 75.54 77.54 83.24 66.67
@@
 No Adapt 42.51 45.40 25.50 42.15 42.84 41.87 46.77 44.88 16.51 44.12 49.36 5.42 51.33 52.35 58.47 40.63
+RoTTA 43.02 44.94 25.48 42.24 43.26 42.13 46.93 44.50 16.63 43.83 49.54 5.45 50.99 52.37 58.50 40.65
+DeYO 43.00 45.45 25.64 42.02 42.88 42.47 46.61 44.97 16.79 43.82 49.30 5.45 51.34 52.85 58.45 40.74
 SITA 43.01 44.83 25.48 42.23 43.29 42.11 46.93 44.49 16.65 43.84 49.53 5.56 50.99 52.48 58.71 40.68
@@
 No Adapt 12.43 15.20 8.74 7.58 6.50 14.48 13.59 15.58 5.27 18.06 15.01 1.32 20.53 31.62 31.15 14.47
+RoTTA 12.67 15.34 8.89 7.55 6.56 14.57 13.94 15.75 5.21 18.06 14.82 1.38 21.53 32.21 31.25 14.65
+DeYO 12.69 15.36 8.56 7.39 6.26 14.38 13.75 15.75 5.29 18.12 14.85 1.41 20.93 31.99 31.47 14.55
 SITA 12.71 15.17 8.87 7.56 6.57 14.57 13.94 15.72 5.32 18.05 14.82 1.38 21.45 32.20 31.51 14.66
@@
 SPACE (ours) 16.71 19.44 11.72 9.62 7.31 16.57 15.61 18.66 6.01 21.02 17.53 1.51 21.96 31.89 31.63 16.48
-In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO and SITA.
+In this section, we compare SPACE’s performance with four existing TTA methods.
 To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using SNN-VGG, on CIFAR-10-C with SNN-ResNet11, on ImageNet series with Spike-driven Transformer V3, and on DVS Gesture-C with SNN-ConvLSTM, to demonstrate the broad applicability of our proposed method.
@@
 No Adapt 67.64 41.94 15.12 41.57
+RoTTA 67.34 43.12 14.87 41.78
+DeYO 67.92 41.42 15.42 41.59
+SITA 67.21 42.05 13.92 41.06
 MEMO 67.57 42.15 14.97 41.56
@@
 The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Table 2, highlight the effectiveness of SPACE.
-Key observations include the following: 1) Superior Performance Across Methods: SPACE consistently outperforms No Adapt, SITA, and MEMO in average accuracy across all datasets under both VGG and ResNet architectures.
+Key observations include: 1) Optimal Performance: SPACE consistently outperforms all baselines in average accuracy across all datasets under both VGG and ResNet architectures.
 2) Broad Applicability: SPACE demonstrates effectiveness across different CNN structures, showcasing its generalizability.
@@
 This highlights its robustness in handling complex datasets like Tiny-ImageNet-C.
-4) Limitations of SITA on SNNs: The SNN-ResNet does not include BN layers, rendering SITA inapplicable.
+4) Vs.
+RoTTA: RoTTA’s memory bank cannot be effectively utilized in this single-sample setting, leading to limited performance gains.
+5) Vs.
+SITA: The SNN-ResNet does not include BN layers, rendering SITA inapplicable.
 Furthermore, SITA performs worse than No Adapt, indicating its ineffectiveness for SNNs.
 This may be attributed to the spike-based transmission in SNNs: directly adjusting BN parameters without gradient guidance can disrupt spike information, leading to significant performance degradation.
-5) Comparison with MEMO: While SPACE generally outperforms MEMO and SITA, MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs.
+6) Vs.
+DeYO: DeYO relies on smooth probability outputs for confidence comparison, which is incompatible with SNN’s discrete, temporal spike-based outputs, introducing significant errors.
+Also, DeYO only works normalization layer.
+7) Vs.
+MEMO: MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs.
 82.88% for SPACE) and CIFAR-100-C (59.29% for MEMO vs.
 58.96% for SPACE).
-These minor differences suggest MEMO may have an advantage under certain corruption types, but they do not detract from SPACE’s overall superiority.
+The minor edge cases hint at MEMO’s advantage under certain corruption types, yet they don’t diminish SPACE’s overall superiority.
 Performance Comparison on Spike-driven Transformer V3
 To explore the real-world impact of SPACE, we conducted tests on diverse datasets, including ImageNet-V2, ImageNet-R, and ImageNet-A.
-Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-Event Transformer V3.
-As shown in Figure 3, SPACE consistently outperforms MEMO, achieving a 1.80% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
+Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-driven Transformer V3.
+As shown in Figure 3, SPACE consistently outperforms all baselines, achieving at least 1.59% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
 Performance Comparison on ConvLSTM-based SNN Model
-SNNs are inherently well-suited for processing event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency.
-The DVS-Gesture dataset [1] provides an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures.
+SNNs are well-suited to event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency.
+The DVS-Gesture dataset [1] is an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures.
 To further explore potential network architectures, we incorporated ConvLSTM.
-As shown in Table 2, our approach achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
+BN-dependent SITA and DeYO is not suitable for this architecture.
+As shown in Table 3, SPACE achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
 Table 3: Performance comparison on DVS Gesture-C with ConvLSTM-based SNN model regarding Accuracy (%). The bold number indicates the best result.
@@
 No Adapt 53.03 54.55 36.74 31.06 92.05 91.67 59.85
+RoTTA 55.38 56.54 37.50 30.68 91.62 92.05 60.68
 MEMO 56.06 55.30 36.74 30.68 90.91 91.67 60.23
@@
 4.3 Ablation Studies
-(a) Local feature maps based SPACE outperforms global feature maps based SPACE.
-(b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
-Figure 4: Results of ablation study.
+Figure 4: a) Local feature maps based SPACE outperforms global feature maps based SPACE. b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
 Local vs. Global Feature Map Similarity
+Definitions.
+Let “global feature maps" refers to treating the entire feature map of size $C\times H\times W$, as a single entity and measuring similarity on this full representation.
+Let the “local feature map" considers each spatial position in the $H\times W$ dimension independently, which is adopted to calculate the similarity and averaging across the $C$ channels in the main paper.
 We compared the performance of local versus global feature map similarity for our proposed method.
-The results in 4(a) clearly illustrate the advantage of using similarity of local feature maps over global feature maps.
+The results in Figure 4a clearly illustrate the decisive advantage of using similarity of local feature maps over global feature maps.
 While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency.
-As shown in 4(a), the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets.
+As shown in Figure 4a, the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets.
 This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation.
-These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
+These findings underscore the distinct roles of each channel in SNNs and further highlight the practical limitations of comparing entire feature maps.
 Effect of SPACE on Feature Map Similarity
-To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset.
-As shown in 4(b), we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain.
+further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset.
+As shown in Figure 4b, we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain.
 The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process.
-Specifically, the "After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples.
+Specifically, the “After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples.
 This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
-5 Conclusion and Limitation
+Table 4: Comparison of computational cost on CIFAR-10-C with SNN-VGG9 model.
+Method RoTTA DeYO SITA MEMO SPACE
+GFLOPs 160.67 318.07 161.48 161.48 158.11
+Accuracy 67.79% 67.99% 66.41% 69.20% 71.03%
+Analyses of Computational Cost
+To evaluate computational efficiency, we compared the GFLOPs of different baselines with the same augmented batch size, and SPACE achieved the lowest GFLOPs, demonstrating its efficiency (refers to Table 4).
+Regarding the additional cost of optimizing spike-behavior consistency, this step is performed only once per input in single-sample setting, without requiring whole model backpropagation or iterative updates.
+Effect of Augmentation Quantity
+In the main paper, we follow MEMO [56] to determine the number of augmentations $M$.
+To assess $M$’s effect, we evaluated the average accuracy and the evaluation time per test point under the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$.
+As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement.
+For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly.
+Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
+Table 5: Performance comparison of SPACE with different augmentation quantities on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%) and Evaluation Time (seconds) per test point.
+No Adapt Number of Augmentation
+2 4 8 16 32 64 128
+Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
+Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
+5 Conclusion
 In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs.
-By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations.
-Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures, demonstrating robustness to domain shifts.
-However, SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios.
-Future work will explore alternative optimization strategies to expand its practicality for neuromorphic hardware and energy-efficient AI.
-Appendix A Kernel-Based Consistency Regularization
+By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, directly addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations.
+Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series distribution shifted datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures and backbones, demonstrating robust, practical resilience to real-world domain shifts.
+Acknowledgments and Disclosure of Funding
+This work was supported by Hong Kong Innovation and Technology Fund (ITF) MSRP grant ITS/018/22MS.
+Appendix A Discussion
+Selection of Alignment Layer
+We compared aligning the classifier’s feature map with aligning extractor layers used in the main paper.
+Empirically, classifier-level alignment occasionally degrades performance on several corruptions, whereas extractor-level alignment is consistently more stable.
+We attribute this to: 1) classifier features being highly task-specific and high-curvature—small per-sample TTA steps can move decision boundaries and induce catastrophic forgetting [5]; 2) lower signal-to-noise at the classifier under pseudo-label or consistency noise, which amplifies domain-specific biases; and 3) greater adaptation overhead when updating deeper, larger parameter blocks.
+In contrast, channel-wise alignment at mid-level layers promotes augmentation-invariant structure while keeping the classifier intact, offering a better stability–plasticity trade-off with lower compute.
+Potential Information Loss in Equation 4
+We aggregate spikes over time, convert discrete spike counts into continuous probability values via a softmax function, then calculate the similarity.
+This transformation is the intentional discarding of the absolute magnitude of spike counts to focus on the relative importance of neurons within a channel.
+By applying softmax, we normalize away noisy fluctuations in absolute firing rates and create a stable spatial saliency map.
+Aligning these maps forces the model to learn which spatial features are consistently most important, directly encouraging the learning of invariant representations.
+In addition, while our alignment objective operates on this spatially-focused abstraction, it simultaneously preserves and refines the inherent temporal characteristics of the SNN.
+The underlying model remains a fully temporal processor, and the spike counts are the direct result of its complex dynamics.
+By demanding consistency in the outcome (the count-based saliency map), our method implicitly forces the network to stabilize the cause (the spike generation process).
+This creates a powerful regularization pressure that reduces temporal jitter and enhances the robustness of spike timing, all without the computational cost and overfitting risks associated with direct, spike-for-spike temporal alignment.
+Why not cosine or KL similarity
+We compute similarity as a per-channel dot product over softmax-normalized spike counts, which is bounded, symmetric, and well-defined even when raw counts are zero.
+Cosine similarity requires $\ell_{2}$ normalization and is numerically fragile under SNN sparsity: many channels yield zero or near-zero vectors, causing division-by-zero or unstable gradients; adding an $\varepsilon$ fixes numerics but introduces a sensitive hyperparameter and slightly degrades performance in our trials.
+KL divergence, $\mathrm{KL}(\mathbf{P}\|\mathbf{Q})$, is asymmetric and becomes ill-conditioned when entries approach zero (common with peaked softmax under corruptions), leading to exploding losses unless additional smoothing is used—again adding tuning burden.
+A symmetrized variant (e.g., JSD) alleviates this but increased compute without improving over the simple dot product.
+Given our per-sample TTA regime, the dot product on per-channel probability maps offers the best robustness–efficiency trade-off, avoiding normalization pitfalls while retaining discriminative alignment.
+Feasibility of Deployment
+On neuromorphic platforms like SpiNNaker2 [14], the data pre-processing unit and neural network inference module are typically separate.
+Batch data augmentation can be performed in the pre-processing stage, outside the neuromorphic core, before feeding data into the system for inference.
+This design allows augmentation to be handled efficiently without impacting the neuromorphic computing core, making our method feasible on such platforms.
+Limitation
+Proposed SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios.
+Future work will explore alternative optimization strategies to expand its practicality for more neuromorphic hardware and energy-efficient AI.
+Appendix B Kernel-Based Consistency Regularization
+Table 6: Performance comparison of SPACE without/with kernel embedding (k.e.) on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%). The bold number indicates the best result.
+Method Noise Blur Weather Digital Average
+Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
+SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
+SPACE + k.e. 77.85 79.34 69.25 71.42 67.64 72.21 74.61 78.48 52.46 79.66 83.25 23.40 75.42 76.13 82.98 70.94
 To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function.
-The channel-wise feature probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
+The channel-wise feature probability distribution $\mathbf{P}$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
 $k(\mathbf{z},\mathbf{z}^{\prime})=\text{exp}(-\frac{\|\mathbf{z}-\mathbf{z}^{\prime}\|^{2}}{2\sigma^{2}}),$ (6)
 where $\sigma$ is the kernel bandwidth.
-The kernel function $k(\mathbf{z},\mathbf{z}^{\prime})$ implicitly defines a mapping $\phi$: $\mathcal{Z}\xrightarrow{}\mathcal{H}$, such that the inner product in $\mathcal{H}$ is given by the kernel:
+The kernel function $k(\mathbf{z},\mathbf{z}^{\prime})$ implicitly defines a mapping $\phi$: $\mathcal{Z}\xrightarrow{}\mathcal{H}$, such that the inner product in $\mathcal{H}$ can be given by the kernel function:
 $k(\mathbf{z},\mathbf{z}^{\prime})=\langle\phi(\mathbf{z}),\phi(\mathbf{z}^{\prime})\rangle_{\mathcal{H}}.$ (7)
@@
 Using the kernel trick, this can be computed without explicitly constructing $\phi(\cdot)$:
-$\begin{split}\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\frac{1}{D^{2}}\|\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}})-\sum^{D}_{d=1}\phi(\mathbf{z}_{j_{d}})\|^{2}_{\mathcal{H}}.\end{split}$ (11)
-Then we average the MMD values across all channels:
+$\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\frac{1}{D^{2}}\Bigg\|\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}})-\sum^{D}_{d=1}\phi(\mathbf{z}_{j_{d}})\Bigg\|^{2}_{\mathcal{H}}.$ (11)
+Then we average the MMD values across channels:
 $\text{MMD}^{2}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j})).$ (12)
 Finally, we integrate the average MMD values into the original loss function $\mathcal{L}$:
-$\mathcal{L}^{*}(\theta;\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (13)
-To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using SNN-VGG9 model, as shown in Table 3.
-The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical.
-This limited effectiveness can be attributed to several factors.
-First, the distribution differences introduced by augmentations, such as noise, blur, and weather effects, are relatively small, reducing the need for advanced alignment mechanisms like kernel embedding.
-Additionally, the existing similarity measures in SPACE already effectively capture the relationships between augmented samples, leaving little room for further optimization.
-Moreover, the added computational complexity of kernel embedding, especially the cost of kernel calculations, may even slightly hinder performance in scenarios where the original approach is already sufficient.
+$\mathcal{L}^{*}(E_{\theta_{E}};\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (13)
+We evaluated augmenting SPACE with a kernel embedding loss (MMD) on CIFAR-10-C using SNN-VGG9 (Table 6).
+The average accuracy is virtually unchanged and occasionally slightly worse than our base alignment.
+We hypothesize three reasons.
+1) Representation sufficiency: the spike-count feature map, coupled with our linear, channel-wise similarity, already captures the essential distributional alignment across augmentations; mapping to an RKHS adds little signal.
+2) Regime mismatch: our augmentations are semantics-preserving and bounded in severity (e.g., AugMix [19]), inducing small, approximately uni-modal shifts where MMD—often helpful under large or multi-modal discrepancies—provides limited additional benefit.
+3) Estimation and cost: per-sample TTA affords only a small augmented batch, making MMD estimates (and their gradients) higher-variance and bandwidth-sensitive, while adding nontrivial compute.
 These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
-Table 3: Performance comparison of SPACE without/with kernel embedding (k.e.) on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%). The bold number indicates the best result.
+Appendix C Additional Experimental Information
+C.1 Details on Datasets
+In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability.
+They are CIFAR-10-C [^1] https://zenodo.org/records/2535967, CIFAR-100-C [^2] https://zenodo.org/records/3555552, Tiny-ImageNet-C [^3] https://zenodo.org/records/2469796 [18], ImageNet-V2444https://huggingface.co/datasets/vaishaal/ImageNetV2/tree/main [42], ImageNet-R [^5] https://github.com/hendrycks/imagenet-r [20], ImageNet-A [^6] https://github.com/hendrycks/natural-adv-examples [21] and DVS Gesture-C [^7] http://research.ibm.com/dvsgesture/ [23].
+CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
+These three datasets are commonly used benchmarks for evaluating the robustness of models under various types of input corruptions.
+They are derived from the original CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets by applying 15 corruption types (e.g., noise, blur, weather, and digital distortions) at 5 severity levels to the original CIFAR-10, CIFAR-100 [31], and Tiny-ImageNet [7] datasets, respectively.
+CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes, each with 50,000 training images and 10,000 testing images.
+Tiny-ImageNet-C, derived from Tiny-ImageNet, includes 200 classes with 100,000 training images and 10,000 validation images.
+ImageNet-V2, ImageNet-R and ImageNet-A
+These three datasets are designed to evaluate model robustness under different real-world challenges.
+ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [7], serving as a test of distribution shift.
+ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes.
+ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses.
+Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
+DVS Gesture-C
+DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset.
+This variant introduces six corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise.
+These corruptions, implemented via the Tonic API [35], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
+C.2 Details on Pretrained Models
+SNN-VGG with BNTT
+ [^8] https://github.com/Intelligent-Computing-Lab-Yale/BNTT-Batch-Normalization-Through-Time
+We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [29] as our main backbone model.
+BNTT decouples parameters along the time axis within each layer, effectively capturing spikes’ temporal dynamics.
+The temporally evolving parameters in BNTT enable neurons to regulate spike rates across time steps, facilitating low-latency, low-energy training from scratch.
+The pre-trained model achieves accuracies of 90.61%, 69.50%, and 58.36% on the CIFAR-10, CIFAR-100, and Tiny-ImageNet test sets, respectively.
+SNN-ResNet
+ [^9] https://github.com/chan8972/Enabling_Spikebased_Backpropagation
+To explore other CNN architecture, we use SNN-ResNet11 [32], which trains SNNs directly via surrogate gradient.
+The pre-trained model used in our experiments can achieve 90.95% accuracy on the original CIFAR-10 test sets.
+Spike-driven Transformer V3
+ [^10] https://github.com/BICLab/Spike-Driven-Transformer-V3
+To evaluate the effectiveness of our work on more complex networks and datasets, we utilized a 19M parameter model in Spike-driven Transformer V3 [52], which achieves 79.60% accuracy on the ImageNet-1K validation set.
+This model has a patch size of 14 $\times$ 14, with an embedding dimension of 360 output by the final feature extractor.
+Instead of using the class token, the entire embedding feature map is employed as the input to the classifier.
+This design naturally aligns with our method, where the channel dimension $C$ corresponds to the embedding dimension, and the spatial dimension $D$ matches the patch size.
+SNN-ConvLSTM
+DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures.
+To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64.
+This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
+C.3 Details on Implementation
+To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input.
+In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings.
+For each set of experiments, we employ the SGD optimizer and evaluate $\eta$ ranging from 0.001 to 0.8, selecting the optimal value based on performance.
+We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.
+Appendix D Additional Ablation Studies
+Table 7: Performance comparison of SPACE with different augmentation strength on CIFAR-10-C (Gaussian Noise) with SNN-VGG9 model regarding Accuracy (%).
+Strength $s$ 1 3 5 7 10 Larger Boundary
+Accuracy 77.98 77.88 77.48 77.41 77.02 74.52
+Effect of Augmentation Strength
+We study the impact of AugMix [19] strength $s$ on SPACE.
+Following MEMO, our default is s=1s{=}1; we sweep $s\in[1,10]$ while keeping all other AugMix settings fixed (mixture width/depth, operators) and respecting AugMix’s semantic bounds (e.g., rotation $\leq 30^{\circ}$, translation $\leq\pm 16$ px).
+As shown in Table 7, accuracy remains stable within these bounds, with only minor declines as $s$ increases.
+When we deliberately exceed the bounds (e.g., rotation $50^{\circ}$, translation $\pm 18$ px), accuracy drops noticeably, indicating that overly aggressive transformations inject harmful noise and destabilize per-sample alignment under the augmented batches.
+In this regime, the alignment objective tends to match augmentation-induced artifacts (noise) rather than invariant, class-relevant structure in the feature maps, thereby undermining TTA and diminishing its gains.
+These results are consistent with AugMix’s design goal of bounded, semantics-preserving perturbations; we therefore adopt s=1s{=}1 by default and recommend staying within AugMix’s prescribed ranges for robust adaptation.
+Effect of Encoding Methods for SNNs
+We assess sensitivity to spike encoding by replacing Poisson rate coding with temporal coding in SNN-VGG9 while keeping architecture, training, and TTA hyperparameters fixed.
+As reported in Figure 5, the temporal-coded baseline is markedly less robust under domain shift than its rate-coded counterpart.
+This aligns with intuition: temporal coding relies on precise spike times and is therefore vulnerable to perturbation-induced timing jitter, whereas rate coding averages over spikes, providing redundancy and noise tolerance.
+Figure 5: Performance of SNNs with different encoding under domain shift. Values are Top-1 accuracy (%); Acc. Loss = Clean - Shifted.
+Encoding Method Clean Shifted Acc. Loss
+Temporal Coding 89.27 58.11 31.16
+Rate Coding 90.61 66.57 24.04
+Table 8 shows that SPACE improves both models, with consistently larger relative gains for the temporal-coded network.
+The mechanism is a temporal-boundary effect: the total spike count is a sensitive proxy for timing stability because small perturbations can shift spikes across the fixed processing window; for spikes near the boundary, tiny jitters flip their contribution from 1 to 0, causing discrete changes in the count.
+By enforcing consistency of total counts across augmentations, SPACE penalizes these boundary-crossing events and implicitly forces the model to stabilizes spike timing without brittle spike-to-spike matching.
+The resulting gradients are stronger in temporal-coded models than in rate-coded ones, yielding larger robustness gains.
+Table 8: Performance comparison of spike encoding methods for SNN-VGG9 model on CIFAR-10-C regarding Accuracy (%). The bold number indicates the performance improvement.
 Method Noise Blur Weather Digital Average
 Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+Temporal Coding
+No Adapt 39.74 39.32 68.42 59.59 54.43 60.19 65.90 69.78 46.21 66.77 78.14 27.26 70.29 46.00 79.59 58.11
+SPACE 45.19 44.00 71.82 64.16 56.31 64.90 68.65 70.29 56.90 72.40 79.09 68.64 70.66 50.06 80.07 64.21 (+6.10)
+Rate Coding
 No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
-SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
-SPACE + k.e. 77.85 79.34 69.25 71.42 67.64 72.21 74.61 78.48 52.46 79.66 83.25 23.40 75.42 76.13 82.98 70.94
-Appendix B Exploration of Alternative Alignment Objectives
-B.1 Two Alternative Alignment Objectives
-Average Membrane Potential (SPACE + a.m.p.)
-In the main paper, we use feature maps based on the spike counts over the entire time window as the alignment objective.
-In this section, we explore an alternative biologically inspired metric: the average membrane potential obtained from the LIF neurons in the final layer of the feature extractor.
-Specifically, for a given test point $\mathbf{x}_{i}$, the membrane potentials are captured in the feature map $\mathbf{U}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times D}$, where $T$, $C$, and $D$ represent the time steps, the number of channels, and the spatial dimensions, respectively.
-For LIF neurons that spike at a specific time step, we use the membrane potential after the LIF reset, which indirectly captures the spiking information.
-To obtain the primary objective feature map $\mathbf{F}(\mathbf{x}_{i})$, we compute the temporal average of $\mathbf{U}(\mathbf{x}_{i})$ along the first dimension.
-The resulting feature map is then used to calculate the loss following Equations 4 and Equation 5 in the main paper.
-Finally, the model is updated via SGD.
-Spikes through All Time Steps (SPACE + s.t.t.)
-Instead of summing the spike counts over the temporal dimension, we explicitly retain the temporal information, resulting in a feature map with an additional dimension that captures the spiking activity at each time step.
-This approach preserves the precise temporal dynamics of the spikes and leverages them directly for alignment.
-Specifically, we align the channel-wise spiking activity across all time steps between each pair of augmented samples.
-The update rules remain consistent with those described in the main paper.
-B.2 Experimental Results and Analysis
+SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03 (+4.46)
 Table 9: Performance comparison of SPACE with different objectives on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%).
@@
 SPACE + s.t.t. 72.74 73.84 57.86 62.91 63.66 64.33 70.94 76.35 43.16 74.77 82.30 21.20 74.24 69.35 83.97 66.11
-We implement the two approaches above and evaluate on CIFAR-10-C dataset with SNN-VGG9 model. The results are shown in Table 4.
-Compared to No Adapt, SPACE + a.m.p.
-improves the generalization ability of the model.
-However, its performance is noticeably inferior to SPACE on the majority of corruptions.
-This may be due to the fact that the membrane potentials of the same sample under different augmentations tend to exhibit similar patterns, making them less effective in capturing the temporal distribution characteristics of SNNs.
-For example, in extreme cases where the membrane potential of a LIF neuron stays near the firing threshold throughout all time steps without crossing it, SPACE + a.m.p.
-fails to capture this information.
-In contrast, SPACE can reflect this behavior through spike counts, which more directly encode spiking activity.
-SPACE + s.t.t.
-fails to improve the generalization ability of the model.
-We hypothesize that this is due to the model overfitting during the alignment process.
-Specifically, the updates made to minimize the alignment loss may inadvertently lead to excessive specialization to the augmented samples, thereby hindering its ability to generalize effectively to unseen corruptions.
-Appendix C Additional Ablation Study
-Effect of Augmentation Quantity per Test Point
-In the main paper, we follow MEMO [47] to determine the number of augmentations $M$.
-To analyze the effect of $M$ on our method, we evaluated the average accuracy and the evaluation time per test point the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$.
-As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement.
-For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly.
-Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
-Table 5: Performance comparison of SPACE with different augmentation quantities on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%) and Evaluation Time (seconds) per test point.
-No Adapt Number of Augmentation
-2 4 8 16 32 64 128
-Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
-Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
-Appendix D Additional Experimental Information
-D.1 More Details on Datasets
-In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability.
-They are CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], ImageNet-A [18] and DVS Gesture-C [20].
-CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
-These three datasets are commonly used benchmarks for evaluating the robustness of models under various types of input corruptions.
-They are derived from the original CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets by applying 15 corruption types (e.g., noise, blur, weather, and digital distortions) at 5 severity levels to the original CIFAR-10, CIFAR-100 [26], and Tiny-ImageNet [6] datasets, respectively.
-CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes, each with 50,000 training images and 10,000 testing images.
-Tiny-ImageNet-C, derived from Tiny-ImageNet, includes 200 classes with 100,000 training images and 10,000 validation images.
-ImageNet V2/R/A
-ImageNet-V2, ImageNet-R, and ImageNet-A are three datasets designed to evaluate model robustness under different real-world challenges.
-ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [6], serving as a test of distribution shift.
-ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes.
-ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses.
-Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
-DVS Gesture-C
-DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset.
-This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise.
-These corruptions, implemented via the Tonic API [29], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
-D.2 More Details on Pretrained Models
-SNN-VGG with BNTT
-We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [24] as our main backbone model.
-BNTT decouples the parameters along the time axis within each layer, effectively capturing the temporal dynamics of spikes.
-The temporally evolving learnable parameters in BNTT enable neurons to regulate their spike rates across different time steps, facilitating low-latency and low-energy training from scratch.
-The pre-trained model achieves accuracies of 90.61%, 69.50% and 58.36% on the original CIFAR-10, CIFAR-100 and Tiny-ImageNet test sets respectively.
-SNN-ResNet
-To explore other CNN architecture, we use SNN-ResNet11 [27], which trains SNNs directly via surrogate gradient.
-The pre-trained model used in our experiments can achieve 90.95% accuracy on the original CIFAR-10 test sets.
-Spike-driven Transformer V3
-To evaluate the effectiveness of our work on more complex networks and datasets, we utilized a 19M parameter model in Spike-driven Transformer V3 [44], which achieves 79.60% accuracy on the ImageNet-1K validation set.
-This model has a patch size of 14 $\times$ 14, with an embedding dimension of 360 output by the final feature extractor.
-Instead of using the class token, the entire embedding feature map is employed as the input to the classifier.
-This design naturally aligns with our method, where the channel dimension $C$ corresponds to the embedding dimension, and the spatial dimension $D$ matches the patch size.
-SNN-ConvLSTM
-DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures.
-To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64, respectively.
-This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
-D.3 More Details on Implementation
-To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input.
-In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings.
-For each set of experiments, we employ the SGD optimizer and evaluate learning rates ranging from 0.001 to 0.8, selecting the optimal value based on performance.
-We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.
+Exploration of Alternative Alignment Objectives
+We evaluate two alternatives to SPACE’s default count-based feature map on CIFAR-10-C with SNN-VGG9, keeping architecture, training, and TTA hyperparameters fixed. Results are summarized in Table 9.
+1) Average Membrane Potential (SPACE + a.m.p.).
+Let $\mathbf{U}(\mathbf{x})\in\mathbb{R}^{T\times C\times D}$ denote the LIF membrane traces from the selected layer (see subsection 3.2).
+For spikes, the post-reset potential is used.
+We form $\mathbf{F}(\mathbf{x})=\tfrac{1}{T}\sum_{t=1}^{T}\mathbf{U}_{t}(\mathbf{x})\in\mathbb{R}^{C\times D}$ and apply the same channel-wise consistency loss as SPACE.
+It improves over No Adapt but is consistently worse than count-based SPACE on most corruptions.
+Membrane traces are smooth and highly correlated across augmentations, making them less sensitive to discrete spiking changes; near-threshold, non-firing trajectories can be indistinguishable in $\mathbf{F}(\mathbf{x})$, whereas counts capture such events directly.
+2) Spikes Through Time (SPACE + s.t.t.).
+Instead of summing over time, we retain the temporal dimension of the spike tensor $\mathbf{S}(\mathbf{x})\in\mathbb{R}^{T\times C\times D}$ and align per-channel activity at each time step across augmentations (loss averaged over $t$); optimization is unchanged.
+This variant does not improve robustness.
+Keeping the full temporal waveform inflates variance and forces the model to match augmentation-specific transients: benign timing jitter (e.g., a spike at t=10t{=}10 vs.
+t=12t{=}12) is penalized as a mismatch, producing conflicting gradients that push spike times to align across views rather than stabilizing the underlying excitability.
+The model thus overfits to view-specific temporal shapes, hurting generalization, while also incurring higher compute and memory.
+We further applied Gaussian smoothing along time before alignment, $\tilde{\mathbf{S}}=g_{\sigma}*\mathbf{S}$, to reduce jitter-induced overfitting.
+Results remained suboptimal.
+With small $\sigma$, smoothing is too weak to suppress shift-induced conflicts; with larger $\sigma$, it erases informative peaks and attenuates counts.
+The core issue is not merely information loss but the learning signal each representation provides: spike counts yield a shift-invariant, macro-level target that is robust to timing jitter and directs clear adjustments to neuron excitability, whereas smoothed sequences still demand matching temporal shapes, preserving shift-sensitivity and inducing noisy, contradictory gradients.
+Consequently, SPACE’s count-based objective offers a more stable and effective alignment signal for our TTA setting.
+Table 10: Wilcoxon signed-rank test on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C with CNN-based SNN models. W denotes the Wilcoxon signed-rank statistic and p is the two-sided p-value.
+Compared Method Metric Noise Blur Weather Digital Average
+Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
+CIFAR-10-C, SNN-VGG9
+vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
+vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
+vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
+vs. MEMO W 15 0 15 15 15 15 15 15 15 15 10 9 0 15 0 15
+p 1.0000 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.4000 0.3333 0.0312 0.0312 1.0000 0.0312
+CIFAR-100-C, SNN-VGG9
+vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. MEMO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 10 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.4000 0.0312
+Tiny-ImageNet-C, SNN-VGG11
+vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. MEMO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+Table 11: Wilcoxon signed-rank test at the per-class level on CIFAR-10-C (Gaussian/Shot Noise) with CNN-based SNN models. W denotes the Wilcoxon signed-rank statistic and p is the two-sided p-value.
+Compared Method Metric Class Index
+0 1 2 3 4 5 6 7 8 9
+Gaussian Noise
+vs. No Adapt W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. RoTTA W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. DeYO W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. SITA W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. MEMO W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+Shot Noise
+vs. No Adapt W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. RoTTA W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. DeYO W 15 15 15 15 15 15 15 15 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
+vs. SITA W 15 15 15 15 15 15 15 14 15 15
+p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0625 0.0312 0.0312
+vs. MEMO W 5 0 0 5 5 0 0 0 14 0
+p 0.7812 1.0000 1.0000 0.7812 0.7812 1.0000 1.0000 1.0000 0.0625 1.0000
+Statistical Significance
+We assess statistical significance using the Wilcoxon signed-rank test.
+Specifically, for each of the three main experiments, we run five independent trials and apply a paired Wilcoxon test between SPACE and each baseline across the five runs; reported $p$-values in Table 10 confirm that SPACE’s gains are statistically significant.
+To further strengthen the evidence, we select Gaussian and Shot Noise from CIFAR-10-C and conduct Wilcoxon signed-rank tests at the per-class level (Table 11), demonstrating that the improvements are consistent across categories rather than driven by a few classes.